### v1

# Short-Range Order Based Ultra Fast Large-Scale Modeling of High-Entropy Alloys

Abstract
High-Entropy Alloys (HEAs) exhibit complex atomic interactions, with short-range order (SRO) playing a critical role in determining their properties. Traditional methods, such as Monte Carlo (MC) and hybrid MC-Molecular Dynamics (MC-MD), are often constrained by system size and computational inefficiency, limiting their applicability for large-scale modeling. We introduce PyHEA, a high-performance Python package with a C++ core, designed to efficiently model HEAs through a combination of global and local search algorithms, incremental SRO calculations, and GPU acceleration. PyHEA delivers an impressive speedup of more than 110,000× over state-of-the-art tools like SCRAPs and ATAT, ensuring accurate SRO calculations with unprecedented efficiency. In practical applications, PyHEA successfully modeled a 256,000-atom Fe-Mn-Cr-Co alloy, replicating literature results and reducing computational time by over tenfold compared to hybrid MC-MD approaches. By eliminating the need for MC-MD hybridization, PyHEA reduces SRO simulation times from hours to minutes, offering a scalable and efficient tool for HEA research. This advancement enables unprecedented scale and efficiency in modeling atomic structures and SRO, accelerating HEA research and development.
Index Terms:
High-Entropy Alloys, Short-Range Order, High Performance Computing, Monte Carlo Search, Molecular Dynamics, GPU acceleration
I Introduction
High-Entropy Alloys (HEAs), introduced by Jien-Wei Yeh and colleagues in 2004 1, represent a transformative class in materials science. Comprising five or more principal elements in near-equimolar ratios (typically 5–35 atomic percent each), HEAs exhibit a suite of remarkable properties, including high strength, hardness, and exceptional resistance to oxidation and corrosion 2, 3, 4, 5. These properties make HEAs promising candidates for applications in extreme environments, such as aerospace, nuclear, and energy sectors, while also challenging foundational theories in alloy design 6, 3, 7, 4.
Recent research has revealed the importance of microstructure, especially short-range order (SRO), in determining HEAs’ unique properties 6, 8. particularly short-range order (SRO), which refers to the local atomic arrangement within an alloy 9, 10, 11, 12, 13. SRO, describing the local atomic arrangement, emerges as a prevalent feature in complex concentrated alloys, even under rapid solidification 14. It significantly impacts mechanical performance and thermal stability, enhancing elastic modulus, stiffness, and, in some cases, ductility through an optimized balance of strength and toughness 15, 13, 6, 16, 17. These findings suggest a potential for precise SRO control in tailoring HEA properties, laying a theoretical foundation for new material development.
However, leveraging SRO in HEA design remains a substantial challenge due to the exponentially large configuration space resulting from multiple principal elements 18, 19, 20. For example, in a 250-atom equiatomic Cantor alloy (CrMnFeCoNi), the configuration space is approximately $\frac{250!}{(50!)^{5}}\approx 10^{170}$ — a complexity that escalates with increasing atom count.
Current methods for modeling SRO in HEAs face scalability limitations 21. Simulation tools like ATAT 22, 23 and SCRAPs24, while effective for small to medium-scale systems, encounter efficiency constraints in larger systems and struggle to capture SRO effects, particularly under low-temperature conditions where SRO becomes critical 25, 26, 27. As system size increases, these methods often become computationally prohibitive, limiting their ability to simulate HEAs’ comprehensive behavior at macroscopic scales.
Hybrid Monte Carlo-Molecular Dynamics (MC-MD) methods have been proposed for simulating larger systems of hundreds of thousands of atoms 28, 18, 29. Although these methods support extensive simulations, they demand significant computational resources and rely on highly accurate interatomic potentials, derived from Density Functional Theory (DFT) or Machine Learning Interatomic Potentials (MLIPs) 30, 31, 32, 33. This dependency further exacerbates computational costs, underscoring the need for an efficient modeling framework capable of handling large HEA systems with specific SRO requirements without prohibitive computational demands.
Figure 1: Influence of input parameters on fitness and execution time. (a) Number of solutions, (b) global iterations, (c) number of tasks, and (d) search depth. The plots illustrate the impact of varying each parameter on the fitness (left axis) and execution time (right axis).
To address these challenges, we introduce PyHEA, a computational framework optimized for large-scale HEA modeling with two core innovations aimed at overcoming scalability and efficiency issues in capturing SRO. First, PyHEA combines global and local search algorithms, which balances broad exploration with detailed refinement, efficiently navigating the expansive configuration space typical in HEAs and significantly reducing modeling times. Second, it leverages high-precision potentials, such as those from DFT and MLIPs, within an incremental SRO calculation framework. This approach allows PyHEA to rapidly evolve SRO values with precision, enabling the construction of accurate, large-scale HEA models within minutes rather than days. By harnessing both robust search techniques and precise local interactions, PyHEA advances HEA modeling capabilities, allowing researchers to conduct simulations at unprecedented scales and detail.
II Results
II-A Parameter Optimization Analysis
Figure 2: Comparison of PyHEA and SCRAPs in terms of fitness accuracy and execution time. (a) and (b) show the distribution of fitness values (lower values indicate higher accuracy) for systems containing 128 atoms and 2000 atoms, respectively. (c) illustrates the best fitness values as a function of system size. (d) presents the execution times for SCRAPs, PyHEA (CPU), and PyHEA (GPU) across various system sizes. Both SCRAPs and PyHEA (CPU) were tested on the same CPU cluster equipped with an Intel Xeon Gold 6132 CPU @ 2.60 GHz using 24 processors, while PyHEA (GPU) utilized an NVIDIA 4090 GPU.
This section details the key input parameters and their roles in the optimization process. These parameters include the number of solutions, global iterations, tasks, and search depth, all of which directly influence the balance between computational cost and the quality of the fitness function, defined here as the degree to which the simulated atomic configurations align with target SRO characteristics.
In this context, the fitness function serves as a metric for evaluating the alignment of atomic configurations with the desired SRO properties, where higher fitness values indicate more optimized structures. As illustrated in Fig. 1, the parameters—comprising the number of solutions, global iterations, tasks, and search depth—significantly affect the trade-off between the quality of fitness and computational efficiency.
As depicted in Fig. 1(a-b), an increase in the number of solutions improves fitness; however, beyond 64 solutions, this enhancement yields diminishing returns while causing a marked rise in execution time. A similar pattern emerges for global iterations, where substantial fitness improvements are observed up to 6 iterations. Beyond this threshold, further gains are minimal, whereas execution time continues to escalate. These results indicate that using 64 to 128 solutions and 6 global iterations achieves an effective balance between precision and computational cost.
Parallelism also plays a critical role in the optimization process. As demonstrated in Fig. 1(c), increasing the number of tasks enhances the consistency of fitness outcomes. Nevertheless, execution time grows significantly after 128 tasks, with only marginal improvements in fitness. Therefore, a task count between 64 and 128 represents an optimal range. Similarly, search depth affects the thoroughness of exploration within each iteration. Although deeper searches (greater than 6) yield better fitness outcomes, they come at the expense of increased computational demands. Based on these findings, a search depth of 4 to 6 strikes an optimal balance between precision and computational efficiency.
Consequently, the recommended parameter set for this optimization process comprises 64-128 solutions, 6 global iterations, 64-128 tasks, and a search depth of 4-6. This configuration ensures scalability and computational feasibility while maintaining high precision in fitness assessments.
Figure 3: Combined visualization of SRO and radial distribution functions (RDF) for atomic pair interactions at different temperatures. (a) The radar plot illustrates the SRO values for key atomic pairs (e.g., Fe-Fe, Fe-Mn, Fe-Cr) at three temperatures: 750 K, 875 K, and 1373 K. (b) The RDF plots display the spatial distribution of atomic pairs as a function of separation distance (Å) for the same temperatures, with smooth curves representing each pair.
II-B Comparison with State-of-the-Art Methods
To assess the performance of PyHEA, we conducted a comparative analysis against SCRAPs 24, a state-of-the-art method commonly utilized in HEA modeling. This comparison focused on evaluating both fitness quality and computational efficiency across varying system sizes.
Figs. 2(a) and 2(b) present the fitness distributions for systems with 128 atoms and 2000 atoms, respectively. PyHEA consistently exhibits a more concentrated fitness distribution compared to SCRAPs, particularly as the system size increases. In the case of the 2000-atom system, PyHEA demonstrates a narrower spread with a higher peak frequency around the optimal fitness value, signifying more stable and reliable outcomes. Conversely, SCRAPs shows a broader distribution of fitness values, indicating that PyHEA’s optimization strategy is more robust, especially in larger systems.
The enhanced fitness quality of PyHEA is further corroborated by Fig. 2(c), which tracks the optimal fitness values as system size increases. While PyHEA and SCRAPs achieve comparable fitness levels for smaller systems (e.g., 128 and 432 atoms), PyHEA surpasses SCRAPs as the system size exceeds 2000 atoms, highlighting its scalability and the effectiveness of its optimization approach. This is particularly crucial for large-scale HEA simulations where fitness convergence is increasingly challenging.
A key advantage of PyHEA is its computational efficiency, particularly with GPU acceleration. Fig. 2(d) compares the execution times of PyHEA (in both CPU and GPU implementations) and SCRAPs as a function of system size. SCRAPs faces significant scalability challenges, with execution times increasing rapidly as the system size grows. Notably, SCRAPs becomes computationally prohibitive at approximately 6750 atoms, beyond which further calculations are unfeasible.
In contrast, PyHEA’s GPU implementation seamlessly handles larger system sizes. At the largest system size of 27,648 atoms, PyHEA’s GPU-accelerated version achieves a performance improvement of four orders of magnitude compared to SCRAPs. This significant boost in computational efficiency allows PyHEA to handle system sizes that are beyond the practical capacity of SCRAPs.
To further validate PyHEA’s acceleration benefits, we also conducted a comparison with another widely used MC-based tool, ATAT. Table I summarizes the performance metrics across PyHEA (CPU and GPU), SCRAPs, and ATAT. While SCRAPs and ATAT perform comparably for smaller system sizes, they both struggle as the number of atoms increases. Like SCRAPs, ATAT becomes impractical beyond 10,000 atoms. Conversely, PyHEA’s GPU implementation maintains substantially lower execution times across all tested system sizes, underscoring its performance advantages and scalability.
II-C Application
TABLE I: PyHEA computation times (in minutes) for generating optimal structures using different hardware configurations, demonstrating speedups over state-of-the-art methods. The search space consisted of 128 solutions, with the global search running for 10 iterations (typically converging within 3 to 5 iterations). Each iteration employed up to 256 parallel MC tasks, each with a search depth of 10. For simplicity, supercell sizes were set as $S=A\cdot L^{3}$, where L denotes the length of the cubic supercell, and A = 2 for body-centered cubic (bcc) or A = 4 for face-centered cubic (fcc), representing the number of atoms in the primitive cell. Although shells can be included to a range that avoids correlation from periodic boundaries, the short-range order (SRO) was optimized over 3 shells. Comparisons are made with popular MC-based ATAT code23 and the SCRAPs method24. Calculations used a cluster (Intel(R) Xeon(R) Gold 6132 CPU @ 2.60GHz/28-threads) and a NVIDIA 4090 GPU.
  Type Species Atoms ATAT SCRAPs serial [24 cores] PyHEA CPU serial [24 cores] PyHEA GPU CPU speedup over SCRAPs [ATAT] GPU speedup over SCRAPs [ATAT]
  bcc 3 54 $>$144024   2.74[ 0.14]  3.22[0.14] 0.028 0.85[447] 98[51429]
bcc 4 128 $>$10,00024   6.51[ 0.34]  4.19[0.19] 0.028  1.55[2387] 233[357143]
bcc 4 432 -  24.22[ 1.26]  5.62[0.28] 0.029  4.31 835
bcc 4 1024 - 103.14[ 3.18]  6.24[0.32] 0.032 16.53 3223
bcc 4 2000 - 276.00[ 7.02]  7.00[0.36] 0.035 39.43 7886
bcc 5 2000 - 285.48[ 7.27]  9.28[0.48] 0.036 30.11 7930
bcc 5 6750 - 639.45[47.44] 14.37[0.71] 0.046 44.50 13901
bcc 6 27648 - - 39.61[2.05] 0.092 - -
  fcc 3 108 -   17.37[ 0.63]   3.89[ 0.17] 0.005  4.47 3474
fcc 4 256 -  68.77[ 2.24]   5.50[ 0.27] 0.006 12.50 11461
fcc 5 500 - 260.16[ 5.81]   9.39[ 0.43] 0.007 27.71 37166
fcc 5 4000 - 1559.2[64.31]  14.17[ 0.65] 0.014 110.0 111370
fcc 4 256000 - - 1167.2[61.08] 0.577 - -
  • a
    Due to excessive computational demands, multinary results cannot be provided.
  • b
    The CPU speedup ratio tests is based on the serial results.
To illustrate the practical potential of PyHEA, we conducted large-scale molecular dynamics (MD) simulations using LAMMPS, modeling a system containing 256,000 atoms of Fe-Mn-Cr-Co (composition: 40% Fe, 40% Mn, 10% Cr, and 10% Co) at various temperatures. The system was generated using PyHEA, enabling us to bypass the conventional Monte Carlo (MC) and MD hybrid process and achieve significant reductions in simulation time.
The system was simulated in an NPT ensemble for 10 ps, with the Radial Distribution Function (RDF) calculated to analyze the SRO behavior at different temperatures. As depicted in Fig. 3, our results align closely with those from a reference study in which MC simulations were performed on a comparable Fe-Mn-Cr-Co system to investigate compositionally-driven SRO34. Additionally, our simulation outcomes were consistent with the ordering patterns reported in experimental studies 35, 36 which highlights the Fe–Mn/Co/Cr preference in the first coordination shell within an fcc $\text{Fe}_{50}\text{Mn}_{30}\text{Cr}_{10}\text{Co}_{10}$ system.
Employing PyHEA to generate input structures for LAMMPS allowed us to circumvent the extensive MC-MD hybrid computations typically required, which can take several days to complete. PyHEA reduced this computation time to just a few minutes, showcasing a remarkable improvement in computational efficiency. Furthermore, the scale of the simulation, involving more than 200,000 atoms, surpasses the capabilities of most existing methods, emphasizing PyHEA’s robustness and effectiveness in handling large-scale atomistic simulations.
III Discussion
This study demonstrates that PyHEA achieves superior accuracy and can efficiently handle significantly larger system sizes compared to traditional methods. While existing tools such as SCRAPs and ATAT show exponential increases in execution time for systems exceeding 6750 atoms and LAMMPS encounters substantial computational bottlenecks, PyHEA’s GPU acceleration supports simulations with over 200,000 atoms (Table I), delivering performance improvements of up to four orders of magnitude. This scalability is critical for investigating material properties at more realistic scales.
A notable achievement of this work is PyHEA’s ability to substantially reduce the time required for SRO simulations. Traditional approaches utilizing LAMMPS or hybrid MC-MD methods often require several days to complete; in contrast, PyHEA’s optimized workflow reduces simulation time to just a few minutes without sacrificing accuracy (Fig. 5). This level of efficiency not only allows for the simulation of extensive system modeling but also ensures accurate capture of complex atomic interactions and SRO.
In practical applications, PyHEA successfully simulated intricate SRO phenomena in a 256,000-atom Fe-Mn-Cr-Co system across various temperatures. The consistency of PyHEA’s RDF results with those reported in existing literature further supports its accuracy and applicability in real-world scenarios. This underscores PyHEA’s value not only as a computational tool but also as a reliable method for conducting HEA simulations. The ability to conduct simulations at this scale opens the door for future research into more complex, multi-component alloy systems, potentially impacting other fields of materials science and alloy design.
PyHEA’s scalability and performance provide broader implications for the research community. The reduction in simulation time from days to minutes facilitates faster research cycles and reduces the computational resources needed, making large-scale atomistic simulations more feasible and cost-effective, even for teams with limited resources. This capability supports more iterative and exploratory research, where rapid adjustments to simulation parameters and re-analysis can lead to deeper insights and faster progress.
In conclusion, PyHEA represents a significant advancement in HEA modeling by addressing the computational challenges associated with large-scale simulations. By integrating global and local search algorithms, GPU optimization, and incremental SRO calculations, PyHEA surpasses state-of-the-art methods such as SCRAPs, ATAT, and LAMMPS in both accuracy and scalability. Its ability to handle systems with over 200,000 atoms while maintaining computational efficiency establishes it as a leading tool for HEA research. The reduction in simulation time, along with consistent fitness evaluations, facilitates a more comprehensive exploration of material properties, particularly the role of SRO in HEA performance.
Future research could focus on further optimizing PyHEA’s algorithms and expanding their application to larger and more complex material systems. Incorporating advanced machine learning models, such as reinforcement learning, could enhance PyHEA’s adaptive capabilities in optimizing SRO calculations. Additionally, collaborative efforts with experimental teams to validate PyHEA’s predictions under different thermomechanical conditions would strengthen its applicability and predictive accuracy. Ultimately, PyHEA signifies a major step forward in HEA modeling, providing a scalable, accurate, and efficient platform for simulating short-range order and atomic interactions in extensive systems.
IV Methods
Figure 4: Workflow of the PyHEA algorithm for optimizing atomic configurations. (a) The global search framework iteratively cycles through the stages of perturbation, local search, and ranking to converge toward an optimal solution. (b) Initial atomic configurations consist of randomly generated and previously identified optimal structures for exploration. (c) Monte Carlo (MC) methods are applied to refine each candidate configuration and evaluate their fitness. (d) Configurations are ranked based on fitness, and the lowest-ranked fraction of solutions is discarded to enhance the selection process for subsequent iterations.
IV-A Global Search
Algorithm 1 Global Search Algorithm
1:Fix input and optimization function
2:Optimized solution
3:Initialize solutions
4:while iteration $<$ Global maximum number do
5:     Create new solutions randomly
6:     Calculate fitness $F$ of the solutions
7:     Choose fraction of solutions with best fitness (top solutions)
8:     for each solution $\in$ top solutions do
9:         Search using Monte Carlo (Local Search)
10:     end for
11:     Discard fraction $p_{a}$ of worst solutions
12:     Rank the solutions and find the current best
13:end while
14:Return the best solutions
As shown in Fig. 4(a), the global search phase is essential for navigating the extensive configuration space of atomic arrangements in HEAs. This phase strikes a balance between exploration and refinement, allowing for the identification of promising regions within the solution space.
The search process begins by generating candidate solutions that represent different atomic configurations (Fig. 4(b)). During each iteration, a subset of solutions undergoes random perturbations to enable broader exploration, while the top-performing configurations are preserved for further refinement (Fig. 4(d)). This approach effectively integrates random exploration with solution improvement.
Drawing inspiration from the Cuckoo Search algorithm, the method employs L’evy flights for optimization 37, 24. These random jumps facilitate global exploration and help prevent the algorithm from becoming trapped in local minima. Each candidate solution is evaluated using a fitness function, $F(X)$, which quantifies the quality of SRO in the atomic structure by comparing local atomic arrangements to the desired SRO characteristics.
To maintain solution diversity, a portion of low-performing solutions is eliminated at the end of each iteration, while the best solutions are carried forward to the next cycle. This iterative procedure continues until convergence, gradually driving the system toward an optimized atomic configuration with enhanced SRO properties.
IV-B Local Search
Algorithm 2 Serial Monte Carlo Search Algorithm
1:Fix input and optimization function
2:Optimized solution
3:Initialize current configuration
4:while iteration $<$ Local iterations do
5:     Perturb configuration randomly
6:     Calculate fitness difference $\delta F=F(x)-F(x_{\text{new}})$
7:     if $\delta F<0$ or $|\delta F|\geq\text{threshold}$ then
8:         Restore fitness and configuration to previous state
9:     end if
10:end while
11:Return the best solutions
Algorithm 3 Parallel Monte Carlo Search Algorithm
1:Fix input and optimization function
2:Optimized solution
3:Initialize current configuration
4:for each task $i$ in parallel do
5:     Load current fitness and configuration
6:     Perturb configuration randomly
7:     Calculate fitness difference $\delta F_{i}=F(x)-F(x_{i})$
8:     Synchronize tasks and sort $\delta F_{i}$ in descending order
9:     if $\delta F_{min}>0$ or $|\delta F_{min}|\leq\text{threshold}$ then
10:         $depth$ = 0
11:         Update current fitness and configuration to $x_{i}$
12:     else
13:         $depth$ += 1
14:     end if
15:     if $depth\geq\text{threshold}$ then
16:         Break
17:     end if
18:end for
19:Return the best solutions
The local search phase plays a critical role in refining candidate solutions obtained from the global search phase, particularly in large-scale HEA simulations. To achieve high performance and precision, a parallelized Monte Carlo (MC) search (Algorithm 3)is employed, leveraging GPU implementation to efficiently refine atomic configurations in parallel. This parallelization enables the distribution of tasks across multiple threads, where configurations are perturbed, fitness differences are evaluated, and only the highest-performing solutions are retained. This method enhances both computational speed and accuracy, significantly reducing the overall computation time.
Compared to the serial version (Algorithm 2), which processes configurations one at a time and struggles with the scalability required for large systems, the parallel approach addresses this limitation by processing multiple configurations simultaneously. This parallelization significantly enhances both scalability and performance. By integrating the parallelized local search with the global search phase, candidate solutions are refined to achieve optimal local configurations, facilitating convergence to high-quality results with improved SRO properties.
IV-C Short-Range Order
Short-Range Order (SRO) characterizes the local atomic structure in HEAs, describing the tendency of certain atomic species to preferentially cluster compared to a random distribution 38. The SRO is typically quantified by the Warren-Cowley parameter 9, denoted as $\alpha_{ij}^{k}$, which measures the degree of ordering between atomic species $i$ and $j$. For a given atom $i$, the SRO parameter for its $k$-th coordination shell is defined as:
$\alpha_{ij}^{k}=1-\frac{P_{ij}^{k}}{c_{j}},$ (1)
where $P_{ij}^{(k)}$ is the conditional probability of finding a $j$ atom in the $k$-th coordination shell of an $i$ atom, $c_{j}$ is the overall concentration of $j$ atoms in the alloy.
The SRO parameter $\alpha_{ij}^{k}$ ranges from $[-1,1]$:
  • $\alpha_{ij}^{k}<0$: Indicates an attraction between elements $i$ and $j$, suggesting a preference to be adjacent.
  • $\alpha_{ij}^{k}>0$: Indicates a repulsion between elements $i$ and $j$, suggesting a preference to be separated.
For each atomic pair type $(i,j)$, the SRO parameter for the entire system of $N$ atoms can be expressed as:
$\langle\alpha_{ij}^{k}\rangle=\frac{1}{N_{i}}\sum_{n=1}^{N_{i}}\left(1-\frac{P_{ij}^{k}}{c_{j}}\right),$ (2)
where $N_{i}$ is the number of atoms of type $i$.
To calculate this parameter, the number of $j$ atoms within each coordination shell around every $i$ atom is determined, and the result is normalized by the total concentration of $j$ atoms ($c_{j}$). computed SRO values for each pair $(i,j)$ are then averaged over the entire system to obtain a global measure of the SRO.
IV-D Molecular Dynamics
Molecular dynamics (MD) simulations were performed using LAMMPS 39, 40, 41, employing the 2-NN MEAM potential for the Fe-Mn-Cr-Co alloy system. The simulations encompassed system sizes ranging from 256 to 260,000 atoms, with a timestep of 2 fs in an NPT ensemble. The built-in MC method in LAMMPS was utilized to compute the Radial Distribution Function (RDF), providing detailed insight into the atomic arrangement and SRO.
These simulation parameters were chosen to ensure reproducibility and to facilitate replication in future studies. The results of these MD simulations are discussed in detail in Section II, demonstrating the effectiveness of our method.
IV-E Incremental Calculation of SRO (Fitness)
Figure 5: Comparison of execution times for incremental and non-incremental fitness and total calculations in Monte Carlo search. The bar chart shows the execution times (in seconds) for different supercell lengths N and corresponding atom counts. Incremental methods demonstrate substantial improvements, reducing execution time for both fitness evaluation and total computation. The incremental approach achieves up to a 99.6% reduction in total computation time for larger supercells.
In the Monte Carlo (MC) search process, the SRO parameter is crucial for evaluating the fitness of each atomic configuration. As previously discussed, SRO quantifies the local atomic arrangement, guiding the search toward optimal configurations during each MC iteration.
In the traditional non-incremental approach, SRO is recalculated from scratch at every MC iteration (Eqs. 2). This recalculation becomes increasingly computationally expensive as system size grows. Fig. 5 illustrates that for larger systems, the SRO calculation can consume more than 99% of the total computational cost, highlighting the necessity for a more efficient method.
To address this challenge, we implemented an incremental SRO calculation. Instead of recalculating the SRO for the entire system at each iteration, we update only the SRO values influenced by the random swap of two atoms, $i$ and $j$. This approach localizes the update to the environments of the swapped atoms and their neighbors, enabling a more efficient adjustment of the SRO during each MC iteration.
For the incremental update, we first compute the change in SRO for the atoms directly involved in the swap. The change in the SRO value for atom $i$ and its neighbors is given by:
$\Delta\alpha_{ij}^{k}=\left(1-\frac{P_{ij}^{k}\text{(new)}}{c_{j}}\right)-\left(1-\frac{P_{ij}^{k}\text{(old)}}{c_{j}}\right),$ (3)
[…]
Benchmark data shown in Fig. 5 highlights the significant computational burden of non-incremental SRO calculations, which account for 97.4% to 99.6% of the total execution time as system sizes increase from 432 to 3,456 atoms. However, the incremental SRO approach substantially alleviates this burden, reducing total execution times to 71.8% for 432 atoms and stabilizing around 65.5% for larger systems. These findings underscore the incremental method’s ability to enhance computational efficiency and scalability for large-scale HEA simulations.
By adopting incremental SRO updates, we maintain the accuracy of fitness evaluations while significantly reducing the computation time for each MC iteration. This ensures that the MC search process remains efficient and scalable for large-scale HEA simulations.
IV-F GPU Optimization
To further accelerate the MC search and enhance scalability, we implemented a parallel GPU-based optimization. Each thread processes an independent task, and each thread block manages the complete MC search for a single solution. This approach allows the GPU to handle multiple solutions simultaneously, leveraging resources such as registers, shared memory, and low-precision data formats to maximize computational efficiency.
IV-F1 Parallel Processing on the GPU
In this setup, each thread performs the following tasks:
  • SRO Calculation: Threads compute the SRO for their assigned atomic configurations using low-precision formats (FP16, Uint16) to reduce data footprint and memory access latency, while maintaining sufficient accuracy.
  • Lattice Structure Storage: The lattice structure 42 is stored in shared memory to minimize memory access time during each MC iteration. This allows all threads within a block to efficiently access and update atomic positions, enhancing computational speed.derived from the SRO values calculated in registers, is updated following each perturbation:
IV-F2 Iterative Fitness Calculation
Each thread in the block computes the fitness of the perturbed solution during every MC iteration. The fitness, derived from the SRO values calculated in registers, is updated following each perturbation:
$F_{\text{new}}=F_{\text{prev}}+\Delta F_{\text{perturb}},$ (4)
where $\Delta F_{\text{perturb}}$ represents the fitness change from the most recent atomic position perturbation. This value is calculated using the incremental SRO update method.
IV-F3 Fitness Sorting and Selection
After computing the updated fitness in registers, threads within the block collaborate to sort their fitness values using shared memory. The sorting process ensures the rapid identification of the best-performing configuration, which updates the current optimal solution stored in shared memory:
$F_{\text{best}}=\min\left(F_{\text{new}}\right)$ (5)
This guarantees that each block retains the best solution found during the current iteration.
IV-F4 Convergence Check and Termination
The MC search continues iterating until convergence is achieved. Convergence is defined as the point where the fitness of the current best solution shows no significant improvement over a fixed number of steps. Specifically, if $\Delta F_{\text{best}}=0$ for a predefined threshold of iterations, the threads within the block terminate the search. The final result, including the atomic configuration and the associated fitness, is then written back to global GPU memory for subsequent analysis.
This GPU-based approach significantly accelerates the MC search by distributing tasks across multiple threads and blocks, enabling the efficient handling of large-scale systems.
V Acknowledgment
This work was conducted under the supervision of Professor Lifeng Liu, School of Integrated Circuits,, Peking University. The author acknowledges the financial support provided by the National Natural Science Foundation of China (Grant No. 61874006).
VI Code Availability
The code used for this study is available at https://github.com/caimeiniu/pyhea. Access is granted under LGPL-3.0 license.

### v2

@@
 High-Entropy Alloys (HEAs) exhibit complex atomic interactions, with short-range order (SRO) playing a critical role in determining their properties.
-Traditional methods, such as Monte Carlo (MC) and hybrid MC-Molecular Dynamics (MC-MD), are often constrained by system size and computational inefficiency, limiting their applicability for large-scale modeling.
-We introduce PyHEA, a high-performance Python package with a C++ core, designed to efficiently model HEAs through a combination of global and local search algorithms, incremental SRO calculations, and GPU acceleration.
-PyHEA delivers an impressive speedup of more than 110,000× over state-of-the-art tools like SCRAPs and ATAT, ensuring accurate SRO calculations with unprecedented efficiency.
-In practical applications, PyHEA successfully modeled a 256,000-atom Fe-Mn-Cr-Co alloy, replicating literature results and reducing computational time by over tenfold compared to hybrid MC-MD approaches.
-By eliminating the need for MC-MD hybridization, PyHEA reduces SRO simulation times from hours to minutes, offering a scalable and efficient tool for HEA research.
-This advancement enables unprecedented scale and efficiency in modeling atomic structures and SRO, accelerating HEA research and development.
-Index Terms:
+Traditional methods, such as Monte Carlo (MC) based ATAT and SCRAPs, and hybrid Monte Carlo-Molecular Dynamics (MC-MD), are often constrained by system size and computational inefficiency, limiting their applicability for large-scale modeling.
+We introduce PyHEA, a Python based toolkit with a high-performance C++ core, designed to efficiently model HEAs through a combination of global and local search algorithms, incremental SRO calculations, and GPU acceleration.
+PyHEA delivers an impressive speedup of more than 333,000× and 10,000× over state-of-the-art tools ATAT and SCRAPs respectively, ensuring accurate SRO calculations with unprecedented efficiency.
+In practical applications, PyHEA successfully modeled a 256,000-atom Fe-Mn-Cr-Co alloy, replicating literature results and reducing computational time from days to seconds with GPU acceleration, achieving over 1,000× speedup compared to hybrid MC-MD approaches.
+This huge breakthrough in computational efficiency revolutionizes HEA modeling by bridging theoretical predictions and practical applications, opening new horizons for the design of next-generation HEAs.
+keywords:
 High-Entropy Alloys , Short-Range Order , High Performance Computing , Monte Carlo Search , Molecular Dynamics , GPU acceleration
-I Introduction
-High-Entropy Alloys (HEAs), introduced by Jien-Wei Yeh and colleagues in 2004 1, represent a transformative class in materials science.
-Comprising five or more principal elements in near-equimolar ratios (typically 5–35 atomic percent each), HEAs exhibit a suite of remarkable properties, including high strength, hardness, and exceptional resistance to oxidation and corrosion 2, 3, 4, 5.
-These properties make HEAs promising candidates for applications in extreme environments, such as aerospace, nuclear, and energy sectors, while also challenging foundational theories in alloy design 6, 3, 7, 4.
-Recent research has revealed the importance of microstructure, especially short-range order (SRO), in determining HEAs’ unique properties 6, 8.
-particularly short-range order (SRO), which refers to the local atomic arrangement within an alloy 9, 10, 11, 12, 13.
-SRO, describing the local atomic arrangement, emerges as a prevalent feature in complex concentrated alloys, even under rapid solidification 14.
-It significantly impacts mechanical performance and thermal stability, enhancing elastic modulus, stiffness, and, in some cases, ductility through an optimized balance of strength and toughness 15, 13, 6, 16, 17.
-These findings suggest a potential for precise SRO control in tailoring HEA properties, laying a theoretical foundation for new material development.
-However, leveraging SRO in HEA design remains a substantial challenge due to the exponentially large configuration space resulting from multiple principal elements 18, 19, 20.
-For example, in a 250-atom equiatomic Cantor alloy (CrMnFeCoNi), the configuration space is approximately $\frac{250!}{(50!)^{5}}\approx 10^{170}$ — a complexity that escalates with increasing atom count.
-Current methods for modeling SRO in HEAs face scalability limitations 21.
-Simulation tools like ATAT 22, 23 and SCRAPs24, while effective for small to medium-scale systems, encounter efficiency constraints in larger systems and struggle to capture SRO effects, particularly under low-temperature conditions where SRO becomes critical 25, 26, 27.
-As system size increases, these methods often become computationally prohibitive, limiting their ability to simulate HEAs’ comprehensive behavior at macroscopic scales.
-Hybrid Monte Carlo-Molecular Dynamics (MC-MD) methods have been proposed for simulating larger systems of hundreds of thousands of atoms 28, 18, 29.
-Although these methods support extensive simulations, they demand significant computational resources and rely on highly accurate interatomic potentials, derived from Density Functional Theory (DFT) or Machine Learning Interatomic Potentials (MLIPs) 30, 31, 32, 33.
-This dependency further exacerbates computational costs, underscoring the need for an efficient modeling framework capable of handling large HEA systems with specific SRO requirements without prohibitive computational demands.
+[…]
+[1]organization=School of Software and Microelectronics, Peking University, city=Beijing, country=China
+[…]
+[2]organization=School of Integrated Circuits, Peking University, city=Beijing, country=China
+1 Introduction
+High-Entropy Alloys (HEAs), introduced by Jien-Wei Yeh et al.
+in 2004 [1], have emerged as a transformative class in materials science.
+Characterized by the presence of five or more principal elements in quasi-equimolar ratios (typically 3 to 55% atomic percent each), HEAs exhibit a unique combination of remarkable properties, including high strength, hardness and exceptional resistance to oxidation and corrosion [2, 3, 4, 5].
+These attributes position HEAs as promising materials for applications in extreme environments, such as aerospace, nuclear, and energy sectors.
+Furthermore, the unconventional nature of HEAs challenges traditional alloy design theories, which spurred significant research interest in their underlying principles [6, 7].
+Recent advancements have highlighted the critical role of microstructure in determining HEAs’ extraordinary properties, particularly the influence of short-range order (SRO) [8, 9].
+SRO refers to the local atomic arrangement within an alloy, a phenomenon that becomes increasingly prominent in complex concentrated alloys, even under rapid solidification conditions [11, 13].
+The impact of SRO on mechanical performance and thermal stability is profound, influencing elastic modulus, stiffness, and ductility by optimizing the balance between strength and toughness [14, 15, 16].
+These findings underscore the potential for leveraging SRO control to tailor HEA properties, providing a theoretical framework for the development of new high-performance materials.
+The primary challenge in designing HEAs with precise SRO control lies in the vast configurational space.
+For an equiatomic HEA with $N$ total atoms and $M$ different principal elements, the configurational space size $\Omega$ is given by:
+$\Omega=\frac{N!}{\left(\frac{N}{M}!\right)^{M}}.$ (1)
+This expression reveals the astronomical complexity of the problem—even a modest Cantor alloy (CrMnFeCoNi) with just 250 atoms yields approximately $10^{170}$ possible configurations [17, 18].
+Moreover, realistic applications demand simulations of tens or even hundreds of thousands of atoms to fully capture mechanical and structural properties [33], further amplifying this complexity.
+Current approaches to HEA modeling face significant limitations in addressing this challenge.
+Static structure construction methods, such as ATAT [21, 42, 22] and SCRAPs [23], rely on Special Quasirandom Structures (SQS) and hybrid Monte Carlo approaches, respectively, to generate random alloy configurations.
+While effective for handling relatively small supercells (tens to hundreds of atoms), these methods become computationally prohibitive for larger systems [20] and fail to capture critical SRO effects, particularly at lower temperatures where atomic ordering significantly influences material properties [24, 25, 26].
+Hybrid Monte Carlo-Molecular Dynamics (MC-MD) methods offer an alternative approach by incorporating MC steps into extended MD simulations.
+Although hybrid MC-MD can theoretically handle systems with hundreds of thousands of atoms [27, 17, 28], its practical effectiveness depends heavily on the availability of highly accurate interatomic potentials.
+Even when such potentials exist, generating configurations that reliably reflect SRO demands substantial computational resources, often requiring hours or days of calculation time.
+To address these challenges, we introduce PyHEA, a high-performance computational framework that revolutionizes large-scale HEA modeling through three key innovations:
+  • 1.
+    Enhanced accuracy and versatility: PyHEA significantly outperforms existing methods like ATAT and SCRAPs in both modeling efficiency and stability.
+Our benchmarks in Table 1 demonstrate up to 333,000 and 13,900 times speedup compared with ATAT and SCRAPs respectively while maintaining or improving accuracy.
+Unlike current tools that primarily generate random configurations, PyHEA supports user-defined target SRO values, making it uniquely suitable for modeling temperature-dependent ordering effects and other practical scenarios where specific atomic arrangements are desired.
+  • 2.
+    Unprecedented scalability and speed: Through advanced GPU acceleration (CUDA) and optimized search algorithms, PyHEA efficiently handles systems with hundreds of thousands of atoms.
+For example, it successfully modeled a 256,000-atom Fe-Mn-Cr-Co system in under 40 seconds—a task that typically requires hours or days using traditional hybrid MC-MD method.
+The details of the hybrid MC-MD simulations are elaborated in section 4.
+This breakthrough in computational efficiency enables the exploration of mechanical properties and structural characteristics at previously inaccessible scales.
+  • 3.
+    Comprehensive software ecosystem: Built on Python with a high-performance C++ core, PyHEA combines accessibility with computational efficiency.
+Released under the LGPL-3.0 license, it provides an end-to-end solution including structure generation, SRO optimization, and analysis tools.
+The modular design facilitates easy integration with existing workflows and enables rapid customization for specific research needs.
+In the following sections, we detail PyHEA’s methodology, demonstrate its performance advantages through comprehensive benchmarks, and showcase its practical applications in a large scale HEA system.
+Our results establish PyHEA as a transformative tool that bridges the gap between theoretical understanding and practical simulation of large-scale HEA systems, opening new possibilities for materials design and optimization.
+2 Results
+2.1 Parameter Optimization Analysis
 Figure 1: Influence of input parameters on fitness and execution time.
@@
 The plots illustrate the impact of varying each parameter on the fitness (left axis) and execution time (right axis).
-To address these challenges, we introduce PyHEA, a computational framework optimized for large-scale HEA modeling with two core innovations aimed at overcoming scalability and efficiency issues in capturing SRO.
-First, PyHEA combines global and local search algorithms, which balances broad exploration with detailed refinement, efficiently navigating the expansive configuration space typical in HEAs and significantly reducing modeling times.
-Second, it leverages high-precision potentials, such as those from DFT and MLIPs, within an incremental SRO calculation framework.
-This approach allows PyHEA to rapidly evolve SRO values with precision, enabling the construction of accurate, large-scale HEA models within minutes rather than days.
-By harnessing both robust search techniques and precise local interactions, PyHEA advances HEA modeling capabilities, allowing researchers to conduct simulations at unprecedented scales and detail.
-II Results
-II-A Parameter Optimization Analysis
-Figure 2: Comparison of PyHEA and SCRAPs in terms of fitness accuracy and execution time.
-(a) and (b) show the distribution of fitness values (lower values indicate higher accuracy) for systems containing 128 atoms and 2000 atoms, respectively.
-(c) illustrates the best fitness values as a function of system size.
-(d) presents the execution times for SCRAPs, PyHEA (CPU), and PyHEA (GPU) across various system sizes.
-Both SCRAPs and PyHEA (CPU) were tested on the same CPU cluster equipped with an Intel Xeon Gold 6132 CPU @ 2.60 GHz using 24 processors, while PyHEA (GPU) utilized an NVIDIA 4090 GPU.
-This section details the key input parameters and their roles in the optimization process.
-These parameters include the number of solutions, global iterations, tasks, and search depth, all of which directly influence the balance between computational cost and the quality of the fitness function, defined here as the degree to which the simulated atomic configurations align with target SRO characteristics.
-In this context, the fitness function serves as a metric for evaluating the alignment of atomic configurations with the desired SRO properties, where higher fitness values indicate more optimized structures.
-As illustrated in Fig.
-1, the parameters—comprising the number of solutions, global iterations, tasks, and search depth—significantly affect the trade-off between the quality of fitness and computational efficiency.
+In optimizing atomic configurations for HEAs, several key parameters significantly influence both the quality of results and computational efficiency.
+These parameters include: the number of solutions (which determines the population size for exploring different atomic arrangements), global iterations (controlling how many times the entire optimization process is repeated), number of tasks (affecting the degree of parallelization in the search process), and search depth (determining how thoroughly each local configuration is explored).
+Each parameter plays a distinct role in the search algorithm detailed in subsection 4.1 and subsection 4.2.
+Understanding their optimal values is crucial for achieving the best balance between computational efficiency and solution quality.
+To evaluate optimization quality, we employ a fitness function that quantifies how well atomic configurations align with the target Short-Range Order (SRO) characteristics.
+Lower fitness values indicate better alignment with desired SRO properties.
+Our analysis reveals distinct patterns in how each parameter affects both fitness and computational efficiency.
 As depicted in Fig.
@@
 Therefore, a task count between 64 and 128 represents an optimal range.
-Similarly, search depth affects the thoroughness of exploration within each iteration.
-Although deeper searches (greater than 6) yield better fitness outcomes, they come at the expense of increased computational demands.
-Based on these findings, a search depth of 4 to 6 strikes an optimal balance between precision and computational efficiency.
-Consequently, the recommended parameter set for this optimization process comprises 64-128 solutions, 6 global iterations, 64-128 tasks, and a search depth of 4-6.
-This configuration ensures scalability and computational feasibility while maintaining high precision in fitness assessments.
-Figure 3: Combined visualization of SRO and radial distribution functions (RDF) for atomic pair interactions at different temperatures.
-(a) The radar plot illustrates the SRO values for key atomic pairs (e.g., Fe-Fe, Fe-Mn, Fe-Cr) at three temperatures: 750 K, 875 K, and 1373 K.
-(b) The RDF plots display the spatial distribution of atomic pairs as a function of separation distance (Å) for the same temperatures, with smooth curves representing each pair.
-II-B Comparison with State-of-the-Art Methods
-To assess the performance of PyHEA, we conducted a comparative analysis against SCRAPs 24, a state-of-the-art method commonly utilized in HEA modeling.
+Similarly, Fig.
+1(d) shows search depth affects the thoroughness of exploration within each iteration.
+Deeper searches (greater than 8) yield better fitness values, they also come at the expense of increased computational demands.
+Based on these findings, a search depth of 8-10 strikes an optimal balance between precision and computational efficiency.
+Consequently, we recommend two parameter configurations optimized for different priorities.
+The precision-optimized configuration (128 solutions, 10 global iterations, 256 tasks, search depth 10) achieves maximum accuracy in SRO calculations.
+The performance-optimized configuration (64 solutions, 6 global iterations, 64 tasks, search depth 8) maintains good optimization quality while reducing computation time as much as possible.
+Users can select either configuration based on their specific requirements for precision versus computational efficiency.
+2.2 Comparison with State-of-the-Art Methods
+Figure 2: Comparison of PyHEA and SCRAPs in terms of fitness accuracy and execution time.
+(a) and (b) show the distribution of fitness values (lower values indicate higher accuracy) for systems containing 128 atoms and 2000 atoms, respectively.
+(c) illustrates the best fitness values as a function of system size.
+(d) presents the execution times for SCRAPs, PyHEA (CPU), and PyHEA (GPU) across various system sizes.
+Both SCRAPs and PyHEA (CPU) were tested on the same CPU cluster equipped with an Intel Xeon Gold 6132 CPU @ 2.60 GHz using 24 processors, while PyHEA (GPU) utilized an NVIDIA 4090 GPU.
+To evaluate the performance of PyHEA, we conducted a comparative analysis against SCRAPs [23], a state-of-the-art method commonly utilized in HEA modeling.
 This comparison focused on evaluating both fitness quality and computational efficiency across varying system sizes.
-Figs.
-2(a) and 2(b) present the fitness distributions for systems with 128 atoms and 2000 atoms, respectively.
+Fig.
+2(a-b) present the fitness distributions for typical bcc configurations with 128 atoms and 2000 atoms, respectively.
 PyHEA consistently exhibits a more concentrated fitness distribution compared to SCRAPs, particularly as the system size increases.
@@
 This is particularly crucial for large-scale HEA simulations where fitness convergence is increasingly challenging.
-A key advantage of PyHEA is its computational efficiency, particularly with GPU acceleration.
-Fig.
+A key advantage of PyHEA is its computational efficiency, particularly with GPU acceleration.Fig.
 2(d) compares the execution times of PyHEA (in both CPU and GPU implementations) and SCRAPs as a function of system size.
@@
 To further validate PyHEA’s acceleration benefits, we also conducted a comparison with another widely used MC-based tool, ATAT.
-Table I summarizes the performance metrics across PyHEA (CPU and GPU), SCRAPs, and ATAT.
+Table 1 summarizes the performance metrics based on serial execution across PyHEA (CPU and GPU), SCRAPs, and ATAT.
 While SCRAPs and ATAT perform comparably for smaller system sizes, they both struggle as the number of atoms increases.
-Like SCRAPs, ATAT becomes impractical beyond 10,000 atoms.
+Like SCRAPs, ATAT becomes impractical beyond a few hundred atoms.
 Conversely, PyHEA’s GPU implementation maintains substantially lower execution times across all tested system sizes, underscoring its performance advantages and scalability.
-II-C Application
-TABLE I: PyHEA computation times (in minutes) for generating optimal structures using different hardware configurations, demonstrating speedups over state-of-the-art methods.
-The search space consisted of 128 solutions, with the global search running for 10 iterations (typically converging within 3 to 5 iterations).
-Each iteration employed up to 256 parallel MC tasks, each with a search depth of 10.
-For simplicity, supercell sizes were set as $S=A\cdot L^{3}$, where L denotes the length of the cubic supercell, and A = 2 for body-centered cubic (bcc) or A = 4 for face-centered cubic (fcc), representing the number of atoms in the primitive cell.
-Although shells can be included to a range that avoids correlation from periodic boundaries, the short-range order (SRO) was optimized over 3 shells.
-Comparisons are made with popular MC-based ATAT code23 and the SCRAPs method24.
-Calculations used a cluster (Intel(R) Xeon(R) Gold 6132 CPU @ 2.60GHz/28-threads) and a NVIDIA 4090 GPU.
-  Type Species Atoms ATAT SCRAPs serial [24 cores] PyHEA CPU serial [24 cores] PyHEA GPU CPU speedup over SCRAPs [ATAT] GPU speedup over SCRAPs [ATAT]
-  bcc 3 54 $>$144024   2.74[ 0.14]  3.22[0.14] 0.028 0.85[447] 98[51429]
-bcc 4 128 $>$10,00024   6.51[ 0.34]  4.19[0.19] 0.028  1.55[2387] 233[357143]
-bcc 4 432 -  24.22[ 1.26]  5.62[0.28] 0.029  4.31 835
-bcc 4 1024 - 103.14[ 3.18]  6.24[0.32] 0.032 16.53 3223
-bcc 4 2000 - 276.00[ 7.02]  7.00[0.36] 0.035 39.43 7886
-bcc 5 2000 - 285.48[ 7.27]  9.28[0.48] 0.036 30.11 7930
-bcc 5 6750 - 639.45[47.44] 14.37[0.71] 0.046 44.50 13901
-bcc 6 27648 - - 39.61[2.05] 0.092 - -
-  fcc 3 108 -   17.37[ 0.63]   3.89[ 0.17] 0.005  4.47 3474
-fcc 4 256 -  68.77[ 2.24]   5.50[ 0.27] 0.006 12.50 11461
-fcc 5 500 - 260.16[ 5.81]   9.39[ 0.43] 0.007 27.71 37166
-fcc 5 4000 - 1559.2[64.31]  14.17[ 0.65] 0.014 110.0 111370
-fcc 4 256000 - - 1167.2[61.08] 0.577 - -
+For a comprehensive comparison of parallel CPU performance between these tools, readers are referred to Table 2 in the Appendix.
+Table 1: Performance comparison between PyHEA and existing methods (ATAT and SCRAPs) for optimizing atomic structures in high-entropy alloys, focusing on serial execution performance (parallel performance data available in Appendix).
+The table shows computation times (in minutes) and speedup ratios for both body-centered cubic (bcc) and face-centered cubic (fcc) structures across varying system sizes (#At) and numbers of species (#Sp).
+Both PyHEA and SCRAPs were executed in serial mode using a single CPU core (Intel Xeon Gold 6132 @ 2.60GHz), with additional GPU (NVIDIA 4090) results shown for PyHEA.
+The optimizations used 128 candidate solutions over 10 global iterations with 256 parallel MC tasks (search depth 10).
+Supercell sizes followed $S=A\cdot L^{3}$, where $A=2$ for bcc or $A=4$ for fcc structures.
+All calculations optimized short-range order (SRO) over 3 coordination shells.
+Dashes (–) indicate computationally infeasible cases, while speedup values are shown relative to SCRAPs[ATAT in brackets].
+  Type #Sp #At ATAT SCRAPs PyHEA PyHEA Speedup
+Serial Serial GPU CPU GPU
+  bcc 3 54 >1.4ka 2.7 3.2 .03 .9[438] 98[47k]
+bcc 4 128 >10ka 6.5 4.2 .03 1.6[2.4k] 233[333k]
+bcc 4 432 24.2 5.6 .03 4.3 835
+bcc 4 1024 103 6.2 .03 16.5 3.2k
+bcc 4 2048 276 7.0 .04 39.4 7.9k
+bcc 5 2048 286 9.3 .04 30.1 7.9k
+bcc 5 6.8k 639 14.4 .05 44.5 13.9k
+bcc 6 28k 39.6 .09
+fcc 3 108 17.4 3.9 .01 4.5 3.5k
+fcc 4 256 68.8 5.5 .01 12.5 11.5k
+fcc 5 500 260 9.4 .01 27.7 37.2k
+fcc 5 4k 1.6k 14.2 .01 110 111k
+fcc 4 256k 1.2k .58
   • a
-    Due to excessive computational demands, multinary results cannot be provided.
+    Notation: #Sp = number of elemental species.
   • b
-    The CPU speedup ratio tests is based on the serial results.
+    Notation: #At = total number of atoms in the system.
+  • c
+    Speedup ratios: value vs SCRAPs [value vs ATAT in brackets].
+  • d
+    ATAT data from Ref. [23]. All times in minutes.
+2.3 Application
+Figure 3: Combined visualization of SRO and radial distribution functions (RDF) for atomic pair interactions at different temperatures.
+(a) The radar plot illustrates the SRO values for key atomic pairs (e.g., Fe-Fe, Fe-Mn, Fe-Cr) at three temperatures: 750 K, 875 K, and 1373 K.
+(b) The RDF plots display the spatial distribution of atomic pairs as a function of separation distance (Å) for the same temperatures, with smooth curves representing each pair.
 To illustrate the practical potential of PyHEA, we conducted large-scale molecular dynamics (MD) simulations using LAMMPS, modeling a system containing 256,000 atoms of Fe-Mn-Cr-Co (composition: 40% Fe, 40% Mn, 10% Cr, and 10% Co) at various temperatures.
@@
 As depicted in Fig.
-3, our results align closely with those from a reference study in which MC simulations were performed on a comparable Fe-Mn-Cr-Co system to investigate compositionally-driven SRO34.
-Additionally, our simulation outcomes were consistent with the ordering patterns reported in experimental studies 35, 36 which highlights the Fe–Mn/Co/Cr preference in the first coordination shell within an fcc $\text{Fe}_{50}\text{Mn}_{30}\text{Cr}_{10}\text{Co}_{10}$ system.
+3, our results align closely with those from a reference study in which MC simulations were performed on a comparable Fe-Mn-Cr-Co system to investigate compositionally-driven SRO [33].
+Additionally, our simulation outcomes were consistent with the ordering patterns reported in experimental studies [34, 35] in a $\text{Fe}_{50}\text{Mn}_{30}\text{Cr}_{10}\text{Co}_{10}$ system.
 Employing PyHEA to generate input structures for LAMMPS allowed us to circumvent the extensive MC-MD hybrid computations typically required, which can take several days to complete.
-PyHEA reduced this computation time to just a few minutes, showcasing a remarkable improvement in computational efficiency.
-Furthermore, the scale of the simulation, involving more than 200,000 atoms, surpasses the capabilities of most existing methods, emphasizing PyHEA’s robustness and effectiveness in handling large-scale atomistic simulations.
-III Discussion
-This study demonstrates that PyHEA achieves superior accuracy and can efficiently handle significantly larger system sizes compared to traditional methods.
-While existing tools such as SCRAPs and ATAT show exponential increases in execution time for systems exceeding 6750 atoms and LAMMPS encounters substantial computational bottlenecks, PyHEA’s GPU acceleration supports simulations with over 200,000 atoms (Table I), delivering performance improvements of up to four orders of magnitude.
-This scalability is critical for investigating material properties at more realistic scales.
-A notable achievement of this work is PyHEA’s ability to substantially reduce the time required for SRO simulations.
-Traditional approaches utilizing LAMMPS or hybrid MC-MD methods often require several days to complete; in contrast, PyHEA’s optimized workflow reduces simulation time to just a few minutes without sacrificing accuracy (Fig.
-5).
-This level of efficiency not only allows for the simulation of extensive system modeling but also ensures accurate capture of complex atomic interactions and SRO.
-In practical applications, PyHEA successfully simulated intricate SRO phenomena in a 256,000-atom Fe-Mn-Cr-Co system across various temperatures.
-The consistency of PyHEA’s RDF results with those reported in existing literature further supports its accuracy and applicability in real-world scenarios.
-This underscores PyHEA’s value not only as a computational tool but also as a reliable method for conducting HEA simulations.
-The ability to conduct simulations at this scale opens the door for future research into more complex, multi-component alloy systems, potentially impacting other fields of materials science and alloy design.
-PyHEA’s scalability and performance provide broader implications for the research community.
-The reduction in simulation time from days to minutes facilitates faster research cycles and reduces the computational resources needed, making large-scale atomistic simulations more feasible and cost-effective, even for teams with limited resources.
-This capability supports more iterative and exploratory research, where rapid adjustments to simulation parameters and re-analysis can lead to deeper insights and faster progress.
-In conclusion, PyHEA represents a significant advancement in HEA modeling by addressing the computational challenges associated with large-scale simulations.
-By integrating global and local search algorithms, GPU optimization, and incremental SRO calculations, PyHEA surpasses state-of-the-art methods such as SCRAPs, ATAT, and LAMMPS in both accuracy and scalability.
-Its ability to handle systems with over 200,000 atoms while maintaining computational efficiency establishes it as a leading tool for HEA research.
-The reduction in simulation time, along with consistent fitness evaluations, facilitates a more comprehensive exploration of material properties, particularly the role of SRO in HEA performance.
-Future research could focus on further optimizing PyHEA’s algorithms and expanding their application to larger and more complex material systems.
-Incorporating advanced machine learning models, such as reinforcement learning, could enhance PyHEA’s adaptive capabilities in optimizing SRO calculations.
-Additionally, collaborative efforts with experimental teams to validate PyHEA’s predictions under different thermomechanical conditions would strengthen its applicability and predictive accuracy.
-Ultimately, PyHEA signifies a major step forward in HEA modeling, providing a scalable, accurate, and efficient platform for simulating short-range order and atomic interactions in extensive systems.
-IV Methods
+PyHEA reduced this computation time to just tens of seconds, showcasing a remarkable improvement in computational efficiency.
+Furthermore, the scale of the simulation, involving 256,000 atoms, surpasses the capabilities of most existing methods, emphasizing PyHEA’s robustness and effectiveness in handling large-scale atomistic simulations.
+The temperature-dependent evolution of SRO in this system is further illustrated by Fig.
+6 in the Appendix, which presents Warren-Cowley SRO parameter matrices at different temperatures (1373 K, 873 K, and 750 K).
+These heat maps reveal how the atomic ordering patterns shift from a nearly random distribution at high temperatures (1373 K) to more pronounced local ordering at lower temperatures (750 K), providing valuable insight into the system’s structural evolution.
+3 Discussion
+This study demonstrates PyHEA’s transformative impact on high-entropy alloy modeling through three key achievements.
+First, PyHEA achieves unprecedented computational efficiency while maintaining high accuracy.
+Our benchmarks show that PyHEA outperforms traditional tools like ATAT and SCRAPs by several orders of magnitude, delivering speedups of up to 333,000× and 13,900× respectively (Table 1).
+This dramatic improvement in performance does not come at the cost of accuracy—PyHEA consistently produces more stable and accurate results across various system sizes, as evidenced by the tighter fitness distributions shown in Fig.
+2.
+Second, PyHEA successfully addresses the scalability challenge that has long constrained HEA modeling.
+While existing methods become computationally prohibitive beyond a few thousand atoms, PyHEA’s GPU-accelerated framework efficiently handles systems exceeding 256,000 atoms.
+This capability represents a significant breakthrough, enabling simulations at experimentally relevant scales that were previously inaccessible.
+The practical impact of this scalability is demonstrated in our Fe-Mn-Cr-Co system simulation, where PyHEA reduced computation time from days to mere seconds while accurately reproducing temperature-dependent SRO patterns documented in previous studies.
+Third, PyHEA’s innovative approach to SRO calculations through incremental updates and GPU optimization establishes a new standard for computational efficiency in materials modeling.
+The incremental SRO algorithm reduces computational overhead by up to 99.6% (Fig.
+5), while GPU acceleration delivers over 1,000× speedup compared to CPU-based calculations.
+These technical advances make it feasible to conduct large-scale atomistic simulations with unprecedented efficiency, opening new possibilities for materials design and optimization.
+The broader implications of PyHEA’s capabilities extend beyond pure computational advancement.
+The dramatic reduction in simulation time from days to minutes facilitates more iterative and exploratory research approaches.
+Researchers can now rapidly test hypotheses, adjust parameters, and explore diverse atomic configurations that were previously impractical due to computational constraints.
+This acceleration of the research cycle has the potential to significantly accelerate the discovery and optimization of new high-entropy alloys.
+Furthermore, PyHEA’s success in modeling complex SRO phenomena in large-scale systems provides valuable insights for understanding structure-property relationships in HEAs.
+The ability to accurately simulate and analyze SRO patterns across different temperatures and compositions, as demonstrated in our Fe-Mn-Cr-Co case study, offers a powerful tool for predicting and optimizing material properties.
+This capability is particularly crucial given the growing recognition of SRO’s role in determining mechanical and thermal properties of HEAs.
+Future developments could further enhance PyHEA’s capabilities in several directions.
+Integration with advanced machine learning models, particularly reinforcement learning approaches, could optimize the search for desired atomic configurations.
+Expanding the framework to incorporate additional physical properties and interaction models would broaden its applicability across materials science.
+Additionally, collaborative efforts with experimental teams could strengthen the validation of PyHEA’s predictions under various thermomechanical conditions.
+In conclusion, PyHEA represents a significant advancement in computational materials science, bridging the gap between theoretical understanding and practical simulation of large-scale HEA systems.
+Its combination of unprecedented computational efficiency, accurate SRO modeling, and scalability to experimentally relevant sizes establishes a new paradigm for materials design and optimization.
+As the field continues to evolve, PyHEA’s framework provides a robust foundation for future innovations in high-entropy alloy research and development.
+4 Methods
 Figure 4: Workflow of the PyHEA algorithm for optimizing atomic configurations.
@@
 (d) Configurations are ranked based on fitness, and the lowest-ranked fraction of solutions is discarded to enhance the selection process for subsequent iterations.
-IV-A Global Search
+4.1 Global Search
 Algorithm 1 Global Search Algorithm
@@
 14:Return the best solutions
+The optimization of atomic configurations in High-Entropy Alloys (HEAs) presents a complex challenge due to the vast solution space.
+As illustrated by Fig.
+4, our global search framework addresses this challenge through an iterative approach that combines systematic exploration with targeted refinement.
 As shown in Fig.
@@
 This approach effectively integrates random exploration with solution improvement.
-Drawing inspiration from the Cuckoo Search algorithm, the method employs L’evy flights for optimization 37, 24.
+Drawing inspiration from the Cuckoo Search algorithm, the method employs L’evy flights for optimization [36, 23].
 These random jumps facilitate global exploration and help prevent the algorithm from becoming trapped in local minima.
@@
 This iterative procedure continues until convergence, gradually driving the system toward an optimized atomic configuration with enhanced SRO properties.
-IV-B Local Search
+4.2 Local Search
 Algorithm 2 Serial Monte Carlo Search Algorithm
@@
 The local search phase plays a critical role in refining candidate solutions obtained from the global search phase, particularly in large-scale HEA simulations.
-To achieve high performance and precision, a parallelized Monte Carlo (MC) search (Algorithm 3)is employed, leveraging GPU implementation to efficiently refine atomic configurations in parallel.
+To achieve high performance and precision, a parallelized Monte Carlo (MC) search (Alg.
+3)is employed, leveraging GPU implementation to efficiently refine atomic configurations in parallel.
 This parallelization enables the distribution of tasks across multiple threads, where configurations are perturbed, fitness differences are evaluated, and only the highest-performing solutions are retained.
 This method enhances both computational speed and accuracy, significantly reducing the overall computation time.
-Compared to the serial version (Algorithm 2), which processes configurations one at a time and struggles with the scalability required for large systems, the parallel approach addresses this limitation by processing multiple configurations simultaneously.
+Compared to the serial version (Alg.
+2), which processes configurations one at a time and struggles with the scalability required for large systems, the parallel approach addresses this limitation by processing multiple configurations simultaneously.
 This parallelization significantly enhances both scalability and performance.
 By integrating the parallelized local search with the global search phase, candidate solutions are refined to achieve optimal local configurations, facilitating convergence to high-quality results with improved SRO properties.
-IV-C Short-Range Order
-Short-Range Order (SRO) characterizes the local atomic structure in HEAs, describing the tendency of certain atomic species to preferentially cluster compared to a random distribution 38.
-The SRO is typically quantified by the Warren-Cowley parameter 9, denoted as $\alpha_{ij}^{k}$, which measures the degree of ordering between atomic species $i$ and $j$.
+4.3 Short-Range Order
+Short-Range Order (SRO) characterizes the local atomic structure in HEAs, describing the tendency of certain atomic species to preferentially cluster compared to a random distribution [37].
+The SRO is typically quantified by the Warren-Cowley parameter [9], denoted as $\alpha_{ij}^{k}$, which measures the degree of ordering between atomic species $i$ and $j$.
 For a given atom $i$, the SRO parameter for its $k$-th coordination shell is defined as:
@@
 The SRO parameter $\alpha_{ij}^{k}$ ranges from $[-1,1]$:
+  • 1.
     $\alpha_{ij}^{k}<0$: Indicates an attraction between elements $i$ and $j$, suggesting a preference to be adjacent.
+  • 2.
     $\alpha_{ij}^{k}>0$: Indicates a repulsion between elements $i$ and $j$, suggesting a preference to be separated.
@@
 computed SRO values for each pair $(i,j)$ are then averaged over the entire system to obtain a global measure of the SRO.
-IV-D Molecular Dynamics
-Molecular dynamics (MD) simulations were performed using LAMMPS 39, 40, 41, employing the 2-NN MEAM potential for the Fe-Mn-Cr-Co alloy system.
-The simulations encompassed system sizes ranging from 256 to 260,000 atoms, with a timestep of 2 fs in an NPT ensemble.
+4.4 Molecular Dynamics
+Molecular dynamics (MD) simulations were performed using LAMMPS [38, 39, 40], employing the 2-NN MEAM potential for the Fe-Mn-Cr-Co alloy system.
+The simulations encompassed system sizes ranging from 256 to 256,000 atoms, with a timestep of 2 fs in an NPT ensemble.
 The built-in MC method in LAMMPS was utilized to compute the Radial Distribution Function (RDF), providing detailed insight into the atomic arrangement and SRO.
 These simulation parameters were chosen to ensure reproducibility and to facilitate replication in future studies.
-The results of these MD simulations are discussed in detail in Section II, demonstrating the effectiveness of our method.
-IV-E Incremental Calculation of SRO (Fitness)
+The results of these MD simulations are discussed in detail in section 2, demonstrating the effectiveness of our method.
+4.5 Incremental Calculation of SRO (Fitness)
 Figure 5: Comparison of execution times for incremental and non-incremental fitness and total calculations in Monte Carlo search.
@@
 As previously discussed, SRO quantifies the local atomic arrangement, guiding the search toward optimal configurations during each MC iteration.
-In the traditional non-incremental approach, SRO is recalculated from scratch at every MC iteration (Eqs.
-2).
+In the traditional non-incremental approach, SRO is recalculated from scratch at every MC iteration (Equation 3).
 This recalculation becomes increasingly computationally expensive as system size grows.
@@
 This ensures that the MC search process remains efficient and scalable for large-scale HEA simulations.
-IV-F GPU Optimization
+4.6 GPU Optimization
 To further accelerate the MC search and enhance scalability, we implemented a parallel GPU-based optimization.
@@
 This approach allows the GPU to handle multiple solutions simultaneously, leveraging resources such as registers, shared memory, and low-precision data formats to maximize computational efficiency.
-IV-F1 Parallel Processing on the GPU
+4.6.1 Parallel Processing on the GPU
 In this setup, each thread performs the following tasks:
+  • 1.
     SRO Calculation: Threads compute the SRO for their assigned atomic configurations using low-precision formats (FP16, Uint16) to reduce data footprint and memory access latency, while maintaining sufficient accuracy.
-  • Lattice Structure Storage: The lattice structure 42 is stored in shared memory to minimize memory access time during each MC iteration.
+  • 2.
+    Lattice Structure Storage: The lattice structure [41] is stored in shared memory to minimize memory access time during each MC iteration.
 This allows all threads within a block to efficiently access and update atomic positions, enhancing computational speed.derived from the SRO values calculated in registers, is updated following each perturbation:
-IV-F2 Iterative Fitness Calculation
+4.6.2 Iterative Fitness Calculation
 Each thread in the block computes the fitness of the perturbed solution during every MC iteration.
@@
 where $\Delta F_{\text{perturb}}$ represents the fitness change from the most recent atomic position perturbation. This value is calculated using the incremental SRO update method.
-IV-F3 Fitness Sorting and Selection
+4.6.3 Fitness Sorting and Selection
 After computing the updated fitness in registers, threads within the block collaborate to sort their fitness values using shared memory.
@@
 This guarantees that each block retains the best solution found during the current iteration.
-IV-F4 Convergence Check and Termination
+4.6.4 Convergence Check and Termination
 The MC search continues iterating until convergence is achieved.
@@
 This GPU-based approach significantly accelerates the MC search by distributing tasks across multiple threads and blocks, enabling the efficient handling of large-scale systems.
-V Acknowledgment
+5 Acknowledgment
 This work was conducted under the supervision of Professor Lifeng Liu, School of Integrated Circuits,, Peking University.
@@
 61874006).
-VI Code Availability
+6 Code Availability
 The code used for this study is available at https://github.com/caimeiniu/pyhea. Access is granted under LGPL-3.0 license.
+Appendix A Supplementary Data
+Table 2: Performance comparison between PyHEA and existing methods under parallel computing conditions.
+Both PyHEA and SCRAPs utilized 24 CPU cores (Intel Xeon Gold 6132 @ 2.60GHz) for parallel execution, with additional GPU (single NVIDIA 4090) results shown for PyHEA.
+The table presents computation times (in minutes) and speedup ratios for both body-centered cubic (bcc) and face-centered cubic (fcc) structures of varying sizes (#At) and species numbers (#Sp).
+All simulations maintained consistent parameters: 128 candidate solutions, 10 global iterations, 256 MC tasks (depth 10), and supercell sizes following $S=A\cdot L^{3}$ where $A=2$ (bcc) or $A=4$ (fcc).
+Short-range order was optimized over 3 coordination shells.
+Dashes (–) indicate computationally infeasible cases, while speedup values are shown relative to SCRAPs[ATAT in brackets].
+  Type #Spa #Atb ATAT SCRAPs PyHEA PyHEA Speedupc
+Parallel Parallel GPU CPU GPU
+  bcc 3 54 >1.4kd 0.14 0.14 .03 1.0[10k] 4.67[47k]
+bcc 4 128 >10kd 0.34 0.19 .03 1.79[53k] 11.33[333k]
+bcc 4 432 1.26 0.28 .03 4.5 42
+bcc 4 1024 3.18 0.32 .03 9.94 106
+bcc 4 2048 7.02 0.36 .04 19.5 175.5
+bcc 5 2048 7.27 0.48 .04 15.2 181.8
+bcc 5 6.8k 47.4 0.71 .05 66.8 948
+bcc 6 28k 2.05 .09
+fcc 3 108 0.63 0.17 .01 3.71 63
+fcc 4 256 2.24 0.27 .01 8.3 224
+fcc 5 500 5.81 0.43 .01 13.5 581
+fcc 5 4k 64.3 0.65 .01 99.0 6.4k
+fcc 4 256k 61.1 .58
+  • a
+    Notation: #Sp = number of elemental species.
+  • b
+    Notation: #At = total number of atoms in the system.
+  • c
+    Speedup ratios: value vs SCRAPs [value vs ATAT in brackets].
+  • d
+    ATAT data from Ref. [23]. All times in minutes.
+Figure 6: Warren–Cowley SRO parameter matrices for a multi-principal-element alloy (Fe, Mn, Cr, Co) at three different temperatures: (a) 1373 K, (b) 873 K, and (c) 873 K.
+The color scale ranges from blue (positive values) to red (negative values), indicating a preference for certain atom types to cluster together (positive SRO) or to avoid each other (negative SRO).
+At high temperatures (1373 K), the SRO values are near zero, reflecting a more random distribution, whereas at lower temperatures (873 K), the emergence of pronounced SRO patterns reveals stronger local atomic ordering.

### v3

@@
 High-Entropy Alloys (HEAs) exhibit complex atomic interactions, with short-range order (SRO) playing a critical role in determining their properties.
-Traditional methods, such as Monte Carlo (MC) based ATAT and SCRAPs, and hybrid Monte Carlo-Molecular Dynamics (MC-MD), are often constrained by system size and computational inefficiency, limiting their applicability for large-scale modeling.
-We introduce PyHEA, a Python based toolkit with a high-performance C++ core, designed to efficiently model HEAs through a combination of global and local search algorithms, incremental SRO calculations, and GPU acceleration.
-PyHEA delivers an impressive speedup of more than 333,000× and 10,000× over state-of-the-art tools ATAT and SCRAPs respectively, ensuring accurate SRO calculations with unprecedented efficiency.
-In practical applications, PyHEA successfully modeled a 256,000-atom Fe-Mn-Cr-Co alloy, replicating literature results and reducing computational time from days to seconds with GPU acceleration, achieving over 1,000× speedup compared to hybrid MC-MD approaches.
-This huge breakthrough in computational efficiency revolutionizes HEA modeling by bridging theoretical predictions and practical applications, opening new horizons for the design of next-generation HEAs.
+Traditional methods, such as Monte Carlo generator of Special Quasirandom Structures within the Alloy Theoretic Automated Toolkit (ATAT-mcsqs), Super-Cell Random APproximates (SCRAPs), and hybrid Monte Carlo-Molecular Dynamics (MC-MD)—are often hindered by limited system sizes and high computational costs.
+In response, we introduce PyHEA, a Python-based toolkit with a high-performance C++ core that leverages global and local search algorithms, incremental SRO computations, and GPU acceleration for unprecedented efficiency.
+When constructing random HEAs, PyHEA achieves speedups exceeding 333,000× and 13,900× over ATAT-mcsqs and SCRAPs, respectively, while maintaining high accuracy.
+PyHEA also offers a flexible workflow that allows users to incorporate target SRO values from external simulations (e.g., LAMMPS or density functional theory (DFT)), thereby enabling more realistic and customizable HEA models.
+As a proof of concept, PyHEA successfully replicated literature results for a 256,000-atom Fe–Mn–Cr–Co system within minutes—an order-of-magnitude improvement over hybrid MC–MD approaches.
+This dramatic acceleration opens new possibilities for bridging theoretical insights and practical applications, paving the way for the efficient design of next-generation HEAs.
 keywords:
 High-Entropy Alloys , Short-Range Order , High Performance Computing , Monte Carlo Search , Molecular Dynamics , GPU acceleration
+journal: Computational Materials Science
 […]
-[1]organization=School of Software and Microelectronics, Peking University, city=Beijing, country=China
+[1]organization=School of Software and Microelectronics, Peking University, city=Beijing, postcode=100871, country=China
 […]
-[2]organization=School of Integrated Circuits, Peking University, city=Beijing, country=China
+[2]organization=School of Integrated Circuits, Peking University, city=Beijing, postcode=100871, country=China
+{highlights}
+High Performance: PyHEA delivers unprecedented speedups of 333,000× over ATAT-mcsqs and 13,900× over SCRAPs in random structure generation, achieving superior efficiency without compromising accuracy.
+Large-Scale Compatibility: PyHEA models systems with over 256,000 atoms in seconds, enabling simulations at scales relevant to experimental studies and unlocking new opportunities for high-entropy alloy research.
+GPU Acceleration: PyHEA’s CUDA-based parallel Monte Carlo implementation provides over 1,000× speedup compared to CPU-based calculations, dramatically reducing simulation times for large-scale configurations.
+User-Friendly Python Ecosystem: Built on a Python interface with a high-performance C++ core, PyHEA provides an accessible yet powerful solution for structure generation, SRO analysis, and data visualization, with seamless integration into existing workflows.
 1 Introduction
@@
 in 2004 [1], have emerged as a transformative class in materials science.
-Characterized by the presence of five or more principal elements in quasi-equimolar ratios (typically 3 to 55% atomic percent each), HEAs exhibit a unique combination of remarkable properties, including high strength, hardness and exceptional resistance to oxidation and corrosion [2, 3, 4, 5].
+Characterized by the presence of five or more principal elements in quasi-equimolar ratios (typically 3-55% atomic percent each), HEAs exhibit a unique combination of remarkable properties, including high strength, hardness, and exceptional resistance to oxidation and corrosion [2, 3, 4, 5].
 These attributes position HEAs as promising materials for applications in extreme environments, such as aerospace, nuclear, and energy sectors.
 Furthermore, the unconventional nature of HEAs challenges traditional alloy design theories, which spurred significant research interest in their underlying principles [6, 7].
-Recent advancements have highlighted the critical role of microstructure in determining HEAs’ extraordinary properties, particularly the influence of short-range order (SRO) [8, 9].
+Recent advancements have highlighted the critical role of microstructure in determining HEAs’ extraordinary properties, particularly the influence of short-range order (SRO) [8, 9, 24, 25, 26].
 SRO refers to the local atomic arrangement within an alloy, a phenomenon that becomes increasingly prominent in complex concentrated alloys, even under rapid solidification conditions [11, 13].
@@
 Current approaches to HEA modeling face significant limitations in addressing this challenge.
-Static structure construction methods, such as ATAT [21, 42, 22] and SCRAPs [23], rely on Special Quasirandom Structures (SQS) and hybrid Monte Carlo approaches, respectively, to generate random alloy configurations.
-While effective for handling relatively small supercells (tens to hundreds of atoms), these methods become computationally prohibitive for larger systems [20] and fail to capture critical SRO effects, particularly at lower temperatures where atomic ordering significantly influences material properties [24, 25, 26].
-Hybrid Monte Carlo-Molecular Dynamics (MC-MD) methods offer an alternative approach by incorporating MC steps into extended MD simulations.
+Static structure construction methods like the Monte Carlo generator of Special Quasirandom Structures within the Alloy Theoretic Automated Toolkit (ATAT-mcsqs) [21, 42, 22] and Super-Cell Random APproximates (SCRAPs) [23] rely on Special Quasirandom Structures (SQS) and hybrid Monte Carlo approaches to generate random alloy configurations respectively.
+While these methods are effective for handling relatively small supercells (tens to hundreds of atoms), they become computationally prohibitive when dealing with larger systems.
+[20]
+In parallel, Hybrid Monte Carlo-Molecular Dynamics (MC-MD) methods offer another alternative approach by incorporating MC steps into extended MD simulations.
 Although hybrid MC-MD can theoretically handle systems with hundreds of thousands of atoms [27, 17, 28], its practical effectiveness depends heavily on the availability of highly accurate interatomic potentials.
@@
   • 1.
-    Enhanced accuracy and versatility: PyHEA significantly outperforms existing methods like ATAT and SCRAPs in both modeling efficiency and stability.
-Our benchmarks in Table 1 demonstrate up to 333,000 and 13,900 times speedup compared with ATAT and SCRAPs respectively while maintaining or improving accuracy.
-Unlike current tools that primarily generate random configurations, PyHEA supports user-defined target SRO values, making it uniquely suitable for modeling temperature-dependent ordering effects and other practical scenarios where specific atomic arrangements are desired.
+    Enhanced efficiency: We benchmarked PyHEA against ATAT-mcsqs and SCRAPs by generating a series of random structures.
+As summarized in Table 1, PyHEA achieves speedups of up to $3.33\times 10^{5}$ over ATAT-mcsqs and $1.39\times 10^{4}$ over SCRAPs, all without sacrificing the accuracy of the generated configurations.
+These results highlight PyHEA’s robust numerical stability and superior modeling efficiency, establishing it as a powerful alternative to existing methods.
   • 2.
-    Unprecedented scalability and speed: Through advanced GPU acceleration (CUDA) and optimized search algorithms, PyHEA efficiently handles systems with hundreds of thousands of atoms.
-For example, it successfully modeled a 256,000-atom Fe-Mn-Cr-Co system in under 40 seconds—a task that typically requires hours or days using traditional hybrid MC-MD method.
-The details of the hybrid MC-MD simulations are elaborated in section 4.
-This breakthrough in computational efficiency enables the exploration of mechanical properties and structural characteristics at previously inaccessible scales.
+    Unprecedented scalability: Beyond random structures, PyHEA allows users to specify target SRO values derived from experiments or prior simulations, enabling the modeling of temperature-dependent ordering and other specialized scenarios.
+When such SRO data are unavailable or difficult to obtain, as introduced in  subsection 2.3, we propose an integrated workflow that employs smaller-scale ab initio (DFT) or LAMMPS simulations to compute the desired SRO.
+PyHEA then leverages these inputs to construct large-scale alloys, effectively streamlining conventional MC–MD processes.
+For instance, in modeling a 256,000-atom Fe–Mn–Cr–Co system, PyHEA provides a 1000× speedup in configuration generation.
+Even accounting for the time required to prepare target SRO values, an overall improvement of at least an order of magnitude is retained, making large-scale HEA simulations far more tractable.
   • 3.
-    Comprehensive software ecosystem: Built on Python with a high-performance C++ core, PyHEA combines accessibility with computational efficiency.
-Released under the LGPL-3.0 license, it provides an end-to-end solution including structure generation, SRO optimization, and analysis tools.
+    User-friendly software ecosystem: Built on Python with a high performance C++ core, PyHEA combines accessibility with computational efficiency.
+Released under the LGPL-3.0 license, it provides an end-to-end solution including structure generation, SRO analysis, and data visualization.
 The modular design facilitates easy integration with existing workflows and enables rapid customization for specific research needs.
@@
 2.1 Parameter Optimization Analysis
-Figure 1: Influence of input parameters on fitness and execution time.
-(a) Number of solutions, (b) global iterations, (c) number of tasks, and (d) search depth.
-The plots illustrate the impact of varying each parameter on the fitness (left axis) and execution time (right axis).
-In optimizing atomic configurations for HEAs, several key parameters significantly influence both the quality of results and computational efficiency.
-These parameters include: the number of solutions (which determines the population size for exploring different atomic arrangements), global iterations (controlling how many times the entire optimization process is repeated), number of tasks (affecting the degree of parallelization in the search process), and search depth (determining how thoroughly each local configuration is explored).
-Each parameter plays a distinct role in the search algorithm detailed in subsection 4.1 and subsection 4.2.
-Understanding their optimal values is crucial for achieving the best balance between computational efficiency and solution quality.
-To evaluate optimization quality, we employ a fitness function that quantifies how well atomic configurations align with the target Short-Range Order (SRO) characteristics.
-Lower fitness values indicate better alignment with desired SRO properties.
-Our analysis reveals distinct patterns in how each parameter affects both fitness and computational efficiency.
-As depicted in Fig.
-1(a-b), an increase in the number of solutions improves fitness; however, beyond 64 solutions, this enhancement yields diminishing returns while causing a marked rise in execution time.
-A similar pattern emerges for global iterations, where substantial fitness improvements are observed up to 6 iterations.
-Beyond this threshold, further gains are minimal, whereas execution time continues to escalate.
-These results indicate that using 64 to 128 solutions and 6 global iterations achieves an effective balance between precision and computational cost.
-Parallelism also plays a critical role in the optimization process.
-As demonstrated in Fig.
-1(c), increasing the number of tasks enhances the consistency of fitness outcomes.
-Nevertheless, execution time grows significantly after 128 tasks, with only marginal improvements in fitness.
-Therefore, a task count between 64 and 128 represents an optimal range.
-Similarly, Fig.
-1(d) shows search depth affects the thoroughness of exploration within each iteration.
-Deeper searches (greater than 8) yield better fitness values, they also come at the expense of increased computational demands.
-Based on these findings, a search depth of 8-10 strikes an optimal balance between precision and computational efficiency.
-Consequently, we recommend two parameter configurations optimized for different priorities.
-The precision-optimized configuration (128 solutions, 10 global iterations, 256 tasks, search depth 10) achieves maximum accuracy in SRO calculations.
-The performance-optimized configuration (64 solutions, 6 global iterations, 64 tasks, search depth 8) maintains good optimization quality while reducing computation time as much as possible.
-Users can select either configuration based on their specific requirements for precision versus computational efficiency.
+Figure 1: Influence of input parameters on fitness (left axis) and execution time (right axis): (a) Number of solutions, (b) global iterations, (c) number of tasks, and (d) search depth.
+In this subsection, we examine how four key parameters influence both the optimization quality and computational efficiency when constructing atomic configurations for HEAs: 1) Number of solutions, 2) Global iterations, 3) Number of tasks, and 4) Search depth.
+These parameters govern the search algorithm outlined in subsection 4.1 and subsection 4.2, while the fitness function used to evaluate candidate configurations is defined in subsection 4.4 (see Equation 3).
+Conceptually, the fitness function serves as a “loss function” in an optimization context: it quantifies how “far” a candidate atomic configuration is from the target SRO.
+Lower fitness values indicate closer alignment with the desired SRO.
+By iteratively minimizing this fitness, the algorithm refines candidate configurations to achieve the prescribed atomic ordering.
+The first parameter, number of solutions, refers to the total number of candidate configurations (atomic arrangements) evaluated in parallel during each global search iteration.
+As illustrated in Fig.
+1(a), increasing the number of solutions from a smaller baseline up to around 64 significantly reduces the fitness, indicating better alignment with the target SRO.
+Beyond 64 solutions, additional gains in accuracy become negligible, while execution time increases sharply.
+Therefore, a range of 64–128 solutions strikes a balance between accuracy and computational efficiency.
+The second parameter, global iterations, specifies the total number of optimization cycles.
+Fig.
+1(b) shows that fitness improves substantially up to about six iterations.
+Beyond this point, further improvements are marginal compared to the additional runtime.
+Hence, six global iterations suffice in most cases to achieve robust SRO alignment without incurring excessive computational costs.
+Parallelism, governed by the number of tasks, is the third parameter under consideration.
+As shown in Fig.
+1(c), increasing the number of tasks up to around 128 enhances the consistency of fitness outcomes.
+However, execution time grows considerably beyond 128 tasks, while fitness gains diminish.
+A task count between 64 and 128 is recommended to maintain an effective balance between throughput and result fidelity.
+The final parameter, search depth, determines the thoroughness of local refinements during each iteration.
+Fig.
+1(d) demonstrates that deeper searches (greater than a depth of 8) lead to lower fitness values, signifying more accurate SRO alignment.
+However, deeper searches also incur higher computational costs.
+Consequently, a search depth of approximately 8–10 provides an optimal trade-off between solution precision and runtime.
+Building on these insights, we propose two configurations tailored to different priorities:
+  • 1.
+    Precision-Optimized Configuration: (128 solutions, 10 global iterations, 256 tasks, search depth 10).
+This setup achieves the highest fidelity to the target SRO but requires longer execution times.
+  • 2.
+    Performance-Optimized Configuration: (64 solutions, 6 global iterations, 64 tasks, search depth 8).
+This arrangement maintains good SRO alignment while minimizing runtime, yielding a practical balance between accuracy and computational efficiency.
+By choosing one of these configurations—or adjusting parameters within the suggested ranges—users can tailor the optimization process to either prioritize maximum accuracy or expedite computations, depending on the demands of their particular HEA simulations.
 2.2 Comparison with State-of-the-Art Methods
@@
 Both SCRAPs and PyHEA (CPU) were tested on the same CPU cluster equipped with an Intel Xeon Gold 6132 CPU @ 2.60 GHz using 24 processors, while PyHEA (GPU) utilized an NVIDIA 4090 GPU.
-To evaluate the performance of PyHEA, we conducted a comparative analysis against SCRAPs [23], a state-of-the-art method commonly utilized in HEA modeling.
+To evaluate the performance of PyHEA, we conducted a comparative analysis against SCRAPs, a state-of-the-art method in HEA modeling[23].
 This comparison focused on evaluating both fitness quality and computational efficiency across varying system sizes.
@@
 This significant boost in computational efficiency allows PyHEA to handle system sizes that are beyond the practical capacity of SCRAPs.
-To further validate PyHEA’s acceleration benefits, we also conducted a comparison with another widely used MC-based tool, ATAT.
-Table 1 summarizes the performance metrics based on serial execution across PyHEA (CPU and GPU), SCRAPs, and ATAT.
-While SCRAPs and ATAT perform comparably for smaller system sizes, they both struggle as the number of atoms increases.
-Like SCRAPs, ATAT becomes impractical beyond a few hundred atoms.
+To further validate PyHEA’s acceleration benefits, we also conducted a comparison with another widely used MC-based tool, ATAT-mcsqs.
+Table 1 summarizes the performance metrics based on serial execution across PyHEA (CPU and GPU), SCRAPs, and ATAT-mcsqs.
+While SCRAPs and ATAT-mcsqs perform comparably for smaller system sizes, they both struggle as the number of atoms increases.
+Like SCRAPs, ATAT-mcsqs becomes impractical beyond a few hundred atoms.
 Conversely, PyHEA’s GPU implementation maintains substantially lower execution times across all tested system sizes, underscoring its performance advantages and scalability.
 For a comprehensive comparison of parallel CPU performance between these tools, readers are referred to Table 2 in the Appendix.
-Table 1: Performance comparison between PyHEA and existing methods (ATAT and SCRAPs) for optimizing atomic structures in high-entropy alloys, focusing on serial execution performance (parallel performance data available in Appendix).
+Table 1: Performance comparison between PyHEA and existing methods (ATAT-mcsqs and SCRAPs) for optimizing atomic structures in HEAs, focusing on serial execution performance (parallel performance data available in Appendix).
 The table shows computation times (in minutes) and speedup ratios for both body-centered cubic (bcc) and face-centered cubic (fcc) structures across varying system sizes (#At) and numbers of species (#Sp).
 Both PyHEA and SCRAPs were executed in serial mode using a single CPU core (Intel Xeon Gold 6132 @ 2.60GHz), with additional GPU (NVIDIA 4090) results shown for PyHEA.
-The optimizations used 128 candidate solutions over 10 global iterations with 256 parallel MC tasks (search depth 10).
+The optimizations used 128 candidate solutions, 10 global iterations, 256 parallel MC tasks, and a search depth of 10.
 Supercell sizes followed $S=A\cdot L^{3}$, where $A=2$ for bcc or $A=4$ for fcc structures.
-All calculations optimized short-range order (SRO) over 3 coordination shells.
+All calculations optimized SRO over 3 coordination shells.
 Dashes (–) indicate computationally infeasible cases, while speedup values are shown relative to SCRAPs[ATAT in brackets].
-  Type #Sp #At ATAT SCRAPs PyHEA PyHEA Speedup
-Serial Serial GPU CPU GPU
+  Type #Sp #At ATAT- SCRAPs PyHEA PyHEA Speedup
+mcsqs Serial Serial GPU CPU GPU
   bcc 3 54 >1.4ka 2.7 3.2 .03 .9[438] 98[47k]
@@
   • d
-    ATAT data from Ref. [23]. All times in minutes.
+    ATAT-mcsqs data from Ref. [23]. All times in minutes.
 2.3 Application
-Figure 3: Combined visualization of SRO and radial distribution functions (RDF) for atomic pair interactions at different temperatures.
-(a) The radar plot illustrates the SRO values for key atomic pairs (e.g., Fe-Fe, Fe-Mn, Fe-Cr) at three temperatures: 750 K, 875 K, and 1373 K.
-(b) The RDF plots display the spatial distribution of atomic pairs as a function of separation distance (Å) for the same temperatures, with smooth curves representing each pair.
-To illustrate the practical potential of PyHEA, we conducted large-scale molecular dynamics (MD) simulations using LAMMPS, modeling a system containing 256,000 atoms of Fe-Mn-Cr-Co (composition: 40% Fe, 40% Mn, 10% Cr, and 10% Co) at various temperatures.
-The system was generated using PyHEA, enabling us to bypass the conventional Monte Carlo (MC) and MD hybrid process and achieve significant reductions in simulation time.
-The system was simulated in an NPT ensemble for 10 ps, with the Radial Distribution Function (RDF) calculated to analyze the SRO behavior at different temperatures.
-As depicted in Fig.
-3, our results align closely with those from a reference study in which MC simulations were performed on a comparable Fe-Mn-Cr-Co system to investigate compositionally-driven SRO [33].
-Additionally, our simulation outcomes were consistent with the ordering patterns reported in experimental studies [34, 35] in a $\text{Fe}_{50}\text{Mn}_{30}\text{Cr}_{10}\text{Co}_{10}$ system.
-Employing PyHEA to generate input structures for LAMMPS allowed us to circumvent the extensive MC-MD hybrid computations typically required, which can take several days to complete.
-PyHEA reduced this computation time to just tens of seconds, showcasing a remarkable improvement in computational efficiency.
-Furthermore, the scale of the simulation, involving 256,000 atoms, surpasses the capabilities of most existing methods, emphasizing PyHEA’s robustness and effectiveness in handling large-scale atomistic simulations.
-The temperature-dependent evolution of SRO in this system is further illustrated by Fig.
-6 in the Appendix, which presents Warren-Cowley SRO parameter matrices at different temperatures (1373 K, 873 K, and 750 K).
-These heat maps reveal how the atomic ordering patterns shift from a nearly random distribution at high temperatures (1373 K) to more pronounced local ordering at lower temperatures (750 K), providing valuable insight into the system’s structural evolution.
+Figure 3: Schematic illustration of the application workflow, with red labels indicating approximate computation times: (1) An initial random structure (containing $\sim 500$ atoms) is generated in under 1 s; (2) this structure is passed to external tools (e.g., LAMMPS or first-principles calculations) for SRO evaluation (taking $\sim 25$ min in LAMMPS); (3) the resulting structure is analyzed in under 1 s to determine the target SRO; (4) a large-scale target structure (containing $256{,}000$ atoms) is constructed in about 35 s, based on the specified SRO parameters.
+As illustrated in Fig.
+3, PyHEA significantly streamlines the modeling workflow for HEAs.
+A small-scale system of approximately 500 atoms is first generated and refined (details of the Monte Carlo steps and convergence criteria are provided in section 4), after which the final SRO parameters are extracted.
+These SRO parameters guide the construction of a large-scale configuration containing up to $256{,}000$ atoms, a process completed in mere 35 seconds—substantially shorter than the hours or days typically required by Monte Carlo–Molecular Dynamics (MC–MD) simulations.
+Figure 4: SRO values and Radial distribution function (RDF) for key atomic pairs at various temperatures in a $256{,}000$-atom Fe–Mn–Cr–Co system.
+(a) A radar plot illustrating SRO values for selected atomic pairs (e.g., Fe–Fe, Fe–Mn, Fe–Cr) at three temperatures: 750 K, 873 K, and 1373 K.
+(b) RDF curves showing the density of atomic pairs as a function of separation distance (Å) at the same temperatures.
+To demonstrate this workflow, we generated a $256{,}000$-atom Fe–Mn–Cr–Co (40 % Fe, 40 % Mn, 10 % Cr, 10 % Co) system using PyHEA, then performed a 10 ps LAMMPS MD simulation in an NPT ensemble at multiple temperatures to obtain radial distribution function (RDF) data.
+Fig.
+4 shows both the RDF and the SRO for representative atomic pairs.
+The RDF peaks indicate the positions of the first and second coordination shells, while the corresponding SRO values reveal how strongly specific atomic species tend to cluster or avoid one another.
+The RDF and SRO curves exhibit a clear transition from a random-like distribution at 1373 K to increasingly pronounced SRO configurations at 750 K.
+This trend aligns qualitatively with prior computational and experimental studies on Fe–Mn–Cr–Co alloys, underscoring the physical reliability and robustness of our approach.
+Finally, additional temperature-dependent SRO details are illustrated in Fig.
+7, which visualizes the Warren–Cowley SRO parameters at 1373 K, 873 K, and 750 K.
+Taken together, these results underscore PyHEA’s efficiency in handling large-scale atomistic simulations and its capability to produce physically meaningful predictions that align well with established literature.
 3 Discussion
-This study demonstrates PyHEA’s transformative impact on high-entropy alloy modeling through three key achievements.
-First, PyHEA achieves unprecedented computational efficiency while maintaining high accuracy.
-Our benchmarks show that PyHEA outperforms traditional tools like ATAT and SCRAPs by several orders of magnitude, delivering speedups of up to 333,000× and 13,900× respectively (Table 1).
-This dramatic improvement in performance does not come at the cost of accuracy—PyHEA consistently produces more stable and accurate results across various system sizes, as evidenced by the tighter fitness distributions shown in Fig.
-2.
-Second, PyHEA successfully addresses the scalability challenge that has long constrained HEA modeling.
-While existing methods become computationally prohibitive beyond a few thousand atoms, PyHEA’s GPU-accelerated framework efficiently handles systems exceeding 256,000 atoms.
-This capability represents a significant breakthrough, enabling simulations at experimentally relevant scales that were previously inaccessible.
-The practical impact of this scalability is demonstrated in our Fe-Mn-Cr-Co system simulation, where PyHEA reduced computation time from days to mere seconds while accurately reproducing temperature-dependent SRO patterns documented in previous studies.
-Third, PyHEA’s innovative approach to SRO calculations through incremental updates and GPU optimization establishes a new standard for computational efficiency in materials modeling.
-The incremental SRO algorithm reduces computational overhead by up to 99.6% (Fig.
-5), while GPU acceleration delivers over 1,000× speedup compared to CPU-based calculations.
-These technical advances make it feasible to conduct large-scale atomistic simulations with unprecedented efficiency, opening new possibilities for materials design and optimization.
-The broader implications of PyHEA’s capabilities extend beyond pure computational advancement.
-The dramatic reduction in simulation time from days to minutes facilitates more iterative and exploratory research approaches.
-Researchers can now rapidly test hypotheses, adjust parameters, and explore diverse atomic configurations that were previously impractical due to computational constraints.
-This acceleration of the research cycle has the potential to significantly accelerate the discovery and optimization of new high-entropy alloys.
-Furthermore, PyHEA’s success in modeling complex SRO phenomena in large-scale systems provides valuable insights for understanding structure-property relationships in HEAs.
-The ability to accurately simulate and analyze SRO patterns across different temperatures and compositions, as demonstrated in our Fe-Mn-Cr-Co case study, offers a powerful tool for predicting and optimizing material properties.
-This capability is particularly crucial given the growing recognition of SRO’s role in determining mechanical and thermal properties of HEAs.
-Future developments could further enhance PyHEA’s capabilities in several directions.
-Integration with advanced machine learning models, particularly reinforcement learning approaches, could optimize the search for desired atomic configurations.
-Expanding the framework to incorporate additional physical properties and interaction models would broaden its applicability across materials science.
-Additionally, collaborative efforts with experimental teams could strengthen the validation of PyHEA’s predictions under various thermomechanical conditions.
-In conclusion, PyHEA represents a significant advancement in computational materials science, bridging the gap between theoretical understanding and practical simulation of large-scale HEA systems.
-Its combination of unprecedented computational efficiency, accurate SRO modeling, and scalability to experimentally relevant sizes establishes a new paradigm for materials design and optimization.
-As the field continues to evolve, PyHEA’s framework provides a robust foundation for future innovations in high-entropy alloy research and development.
+This study presents PyHEA as a computational framework that significantly advances HEA modeling by offering dramatic improvements in simulation speed without compromising accuracy.
+Compared to established tools such as ATAT-mcsqs and SCRAPs, PyHEA achieves speedups of up to 333,000× and 13,900×, respectively (Table 1), while consistently demonstrating tighter fitness distributions (see Fig.
+2).
+These findings confirm that PyHEA maintains or even enhances precision across a wide range of system sizes.
+A key factor in PyHEA’s scalability is its ability to handle systems with more than 256,000 atoms.
+Previous approaches often become intractable beyond a few thousand atoms, yet PyHEA’s GPU-accelerated design and incremental fitness calculations (detailed in subsection 4.5) make large scale simulations both feasible and efficient.
+In the Fe–Mn–Cr–Co case study, for instance, PyHEA reduces computation time from several hours or days to minutes while reliably reproducing the temperature-dependent SRO patterns demonstrated in previous work.
+To explain why PyHEA outperforms existing methods such as SCRAPs, it is crucial to highlight two technical innovations.
+First, the incremental SRO algorithm recalculates only the local environment affected by each atomic swap rather than recomputing the entire system’s SRO at every iteration.
+As described in Section 4.5, this approach lowers SRO-related overhead by up to 99.6%.
+Second, PyHEA capitalizes on GPU acceleration to parallelize local Monte Carlo perturbations and incremental fitness evaluations, thereby multiplying the benefits of the incremental SRO strategy.
+In contrast, SCRAPs relies on a more conventional, system-wide recalculation of SRO and does not exploit large-scale GPU parallelization; hence it cannot match PyHEA’s order-of-magnitude speedups.
+Beyond delivering efficiency and accuracy, PyHEA significantly accelerates the research cycle itself.
+By reducing simulation time from days to mere minutes or seconds, researchers can now quickly test hypotheses, tune parameters, and explore diverse atomic configurations that were previously too time-consuming.
+This capability supports a more iterative approach to materials design, enabling deeper and faster insights into composition–structure–property relationships in HEAs.
+Moreover, PyHEA’s ability to simulate large systems with complex SRO phenomena under various temperatures and compositions provides valuable information for understanding structure-sensitive properties.
+In particular, it opens opportunities for investigating how SRO influences mechanical and thermal behaviors.
+The demonstrated consistency with both experimental and prior computational data reinforces the reliability of PyHEA’s predictions in practical contexts.
+Looking ahead, PyHEA can be extended through integration with machine learning models, especially reinforcement learning, to guide or refine its global search.
+Incorporating more sophisticated interatomic potentials or linking to first-principles simulations would further broaden its applicability.
+Additional validation in collaboration with experimental teams would strengthen confidence in PyHEA’s predictive accuracy under different thermomechanical conditions.
+In conclusion, PyHEA represents a substantial breakthrough in computational materials science by bridging the gap between theory and large-scale HEA simulations.
+Its use of incremental SRO calculations and GPU acceleration explains its marked speedups over methods like SCRAPs.
+By enabling full-atomistic simulations of tens or hundreds of thousands of atoms in a fraction of the time, PyHEA paves the way for faster, more exploratory HEA research and optimization, thus laying a foundation for future advancements in alloy development and discovery.
 4 Methods
+This section details the methods employed in our framework for optimizing atomic configurations in HEAs.
+We discuss the global search at subsection 4.1 for exploring the configuration space, the local search at subsection 4.2 for refining promising candidates, SRO metrics at subsection 4.3 for quantifying atomic arrangements, fitness calculation at subsection 4.4 for evaluating configurations, incremental updates at subsection 4.5 for enhancing computational efficiency, GPU implementation at subsection 4.6 for boosting the performance with heterogeneous computing devices, and finally MD simulations at subsection 4.7 for structure optimization and analysis.
+Fig.
+5 provides an overview of the PyHEA workflow, summarizing its main components and their interactions.
 Figure 5: Workflow of the PyHEA algorithm for optimizing atomic configurations.
@@
 (c) Monte Carlo (MC) methods are applied to refine each candidate configuration and evaluate their fitness.
-(d) Configurations are ranked based on fitness, and the lowest-ranked fraction of solutions is discarded to enhance the selection process for subsequent iterations.
+(d) Configurations are ranked based on fitness, and the lowest-ranked fraction of solutions is discarded to enhance the selection process in subsequent iterations.
 4.1 Global Search
 Algorithm 1 Global Search Algorithm
-1:Fix input and optimization function
-2:Optimized solution
-3:Initialize solutions
+1:Input parameters, optimization function
+2:Optimized solution(s)
+3:Initialize solutions (candidate configurations)
 4:while iteration $<$ Global maximum number do
-5:     Create new solutions randomly
-6:     Calculate fitness $F$ of the solutions
-7:     Choose fraction of solutions with best fitness (top solutions)
-8:     for each solution $\in$ top solutions do
-9:         Search using Monte Carlo (Local Search)
+5:     Create new solutions via random perturbations
+6:     Calculate fitness $F$ for all solutions
+7:     Select a fraction of solutions with the best fitness (top solutions)
+8:     for each solution in top solutions do
+9:         Refine solution via Monte Carlo (local search)
 10:     end for
-11:     Discard fraction $p_{a}$ of worst solutions
-12:     Rank the solutions and find the current best
+11:     Discard the worst fraction $p_{a}$ of solutions
+12:     Rank the solutions and identify the current best
 13:end while
-14:Return the best solutions
-The optimization of atomic configurations in High-Entropy Alloys (HEAs) presents a complex challenge due to the vast solution space.
-As illustrated by Fig.
-4, our global search framework addresses this challenge through an iterative approach that combines systematic exploration with targeted refinement.
-As shown in Fig.
-4(a), the global search phase is essential for navigating the extensive configuration space of atomic arrangements in HEAs.
-This phase strikes a balance between exploration and refinement, allowing for the identification of promising regions within the solution space.
-The search process begins by generating candidate solutions that represent different atomic configurations (Fig.
-4(b)).
-During each iteration, a subset of solutions undergoes random perturbations to enable broader exploration, while the top-performing configurations are preserved for further refinement (Fig.
-4(d)).
-This approach effectively integrates random exploration with solution improvement.
-Drawing inspiration from the Cuckoo Search algorithm, the method employs L’evy flights for optimization [36, 23].
-These random jumps facilitate global exploration and help prevent the algorithm from becoming trapped in local minima.
-Each candidate solution is evaluated using a fitness function, $F(X)$, which quantifies the quality of SRO in the atomic structure by comparing local atomic arrangements to the desired SRO characteristics.
-To maintain solution diversity, a portion of low-performing solutions is eliminated at the end of each iteration, while the best solutions are carried forward to the next cycle.
-This iterative procedure continues until convergence, gradually driving the system toward an optimized atomic configuration with enhanced SRO properties.
+14:return best solutions
+The optimization of atomic configurations in HEAs presents a complex challenge due to the vast solution space.
+As illustrated in Fig.
+5(a), our global search framework implements an iterative optimization strategy that progressively refines atomic configurations through multiple cycles of exploration and evaluation, following the main workflow outlined in Algorithm 1.
+The effectiveness of this strategy is governed by several key parameters, including the number of candidate solutions and global iterations, which significantly influence the balance between exploration breadth and computational efficiency (as detailed in  subsection 2.1).
+The search process begins with the initialization of candidate solutions representing different atomic arrangements (Fig.
+5(b)).
+The number of these initial solutions directly affects the search space coverage, with larger solution pools enabling more comprehensive exploration at the cost of increased computational overhead.
+During the first iteration, these configurations are generated entirely randomly to ensure comprehensive coverage of the solution space.
+In subsequent iterations, the solution pool combines newly generated random configurations with top-performing solutions retained from the previous iteration.
+Crucially, we maintain the number of retained solutions at less than or equal to half of the total configurations selected for local search, thereby ensuring sufficient randomness in each iteration while preventing premature convergence to local optima.
+Each candidate solution undergoes fitness evaluation (subsection 4.4) to quantify its alignment with target SRO characteristics.
+Following evaluation, a predetermined number of the highest-ranking configurations are selected for refinement through local search (Fig.
+5(c)).
+The number of parallel tasks during this phase and the search depth of the Monte Carlo optimization significantly impact both the quality of refinement and computational efficiency.
+This local optimization phase, detailed in  subsection 4.2, employs Monte Carlo perturbations to systematically explore the neighborhood of promising configurations, thereby improving their SRO properties.
+The refined configurations then undergo a comprehensive ranking process, as illustrated by Fig.
+5(d).
+At each iteration’s conclusion, a fixed fraction $p_{a}$ of the lowest-performing solutions is eliminated, while the top-performing configurations are preserved as seeds for the subsequent iteration.
+This selection mechanism, combined with the continuous introduction of new random configurations, maintains an effective balance between preserving successful structural patterns and exploring new possibilities within the vast configuration space.
+This iterative optimization process, as described in Algorithm 1, continues until reaching the maximum number of global iterations.
+The optimal number of iterations, as demonstrated in  subsection 2.1, is crucial for achieving convergence while maintaining computational efficiency.
+Through careful parameter tuning, the algorithm progressively converges toward optimal atomic configurations that exhibit desired SRO characteristics.
 4.2 Local Search
 Algorithm 2 Serial Monte Carlo Search Algorithm
-1:Fix input and optimization function
-2:Optimized solution
-3:Initialize current configuration
-4:while iteration $<$ Local iterations do
-5:     Perturb configuration randomly
-6:     Calculate fitness difference $\delta F=F(x)-F(x_{\text{new}})$
+1:Configuration $x$, optimization function $F$
+2:Refined configuration
+3:Initialize current configuration $x_{\text{curr}}\leftarrow x$
+4:while iteration $<$ Local_iterations do
+5:     Generate a new solution $x_{\text{new}}$
+6:     Calculate $\delta F=F(x_{\text{curr}})-F(x_{\text{new}})$
 7:     if $\delta F<0$ or $|\delta F|\geq\text{threshold}$ then
-8:         Restore fitness and configuration to previous state
+8:         Update $x_{\text{curr}}\leftarrow x_{\text{new}}$
 9:     end if
 10:end while
-11:Return the best solutions
+11:return $x_{\text{curr}}$
 Algorithm 3 Parallel Monte Carlo Search Algorithm
-1:Fix input and optimization function
-2:Optimized solution
-3:Initialize current configuration
+1:Configuration $x$, optimization function $F$
+2:Refined configuration
+3:Initialize current configuration $x_{\text{curr}}\leftarrow x$
 4:for each task $i$ in parallel do
-5:     Load current fitness and configuration
-6:     Perturb configuration randomly
-7:     Calculate fitness difference $\delta F_{i}=F(x)-F(x_{i})$
-8:     Synchronize tasks and sort $\delta F_{i}$ in descending order
-9:     if $\delta F_{min}>0$ or $|\delta F_{min}|\leq\text{threshold}$ then
-10:         $depth$ = 0
-11:         Update current fitness and configuration to $x_{i}$
+5:     Load $x_{\text{curr}}$ and $F(x_{\text{curr}})$
+6:     Generate a new solution $x_{i}$
+7:     Calculate $\delta F_{i}=F(x_{\text{curr}})-F(x_{i})$
+8:     Synchronize tasks, then sort all $\delta F_{i}$ values
+9:     if $\delta F_{\min}>0$ or $|\delta F_{\min}|\leq\text{threshold}$ then
+10:         $depth\leftarrow 0$
+11:         Update $x_{\text{curr}}\leftarrow x_{i}$
 12:     else
-13:         $depth$ += 1
+13:         $depth\leftarrow depth+1$
 14:     end if
@@
 18:end for
-19:Return the best solutions
-The local search phase plays a critical role in refining candidate solutions obtained from the global search phase, particularly in large-scale HEA simulations.
-To achieve high performance and precision, a parallelized Monte Carlo (MC) search (Alg.
-3)is employed, leveraging GPU implementation to efficiently refine atomic configurations in parallel.
-This parallelization enables the distribution of tasks across multiple threads, where configurations are perturbed, fitness differences are evaluated, and only the highest-performing solutions are retained.
-This method enhances both computational speed and accuracy, significantly reducing the overall computation time.
-Compared to the serial version (Alg.
-2), which processes configurations one at a time and struggles with the scalability required for large systems, the parallel approach addresses this limitation by processing multiple configurations simultaneously.
-This parallelization significantly enhances both scalability and performance.
-By integrating the parallelized local search with the global search phase, candidate solutions are refined to achieve optimal local configurations, facilitating convergence to high-quality results with improved SRO properties.
+19:return $x_{\text{curr}}$
+While the global search framework effectively identifies promising configurations through iterative exploration, achieving optimal SRO characteristics requires more focused refinement.
+To address this need, we implement a specialized local search phase employing Monte Carlo techniques with both serial (Algorithm 2) and parallel (Algorithm 3) approaches, with the latter specifically optimized for GPU acceleration to handle large-scale systems efficiently.
+In the serial implementation, each configuration undergoes sequential refinement through iterative solution generation and evaluation.
+The algorithm computes the change in fitness ($\delta F$) for each new solution, accepting modifications that either improve the configuration ($\delta F<0$) or meet specific threshold criteria.
+While this approach is effective for smaller systems, it becomes computationally intensive as system size increases, motivating our development of a parallel implementation.
+The parallel Monte Carlo search (Algorithm 3) significantly enhances efficiency by simultaneously evaluating multiple candidate solutions.
+For each solution, multiple tasks operate in parallel, where each task generates a new configuration by randomly swapping the positions of two atoms.
+As demonstrated in subsection 2.1, the number of parallel tasks represents a crucial parameter, with our analysis revealing optimal performance in the range of 64–128 tasks.
+After each round of atomic swaps, all tasks synchronize to evaluate the fitness changes and identify the most promising configurations.
+This coordinated approach ensures that computational resources are focused on the most promising regions of the solution space, while the depth-based termination criterion prevents excessive computation in regions unlikely to yield significant improvements.
+A key innovation in our methods is the introduction of the search depth parameter, which serves as both a convergence criterion and an efficiency control mechanism.
+As shown in Algorithm 3, the depth counter ($depth$) tracks consecutive non-improving iterations, incrementing when no better solutions are found and resetting to zero when improvements occur.
+Our parametric studies (subsection 2.1) indicate that a search depth threshold of 8–10 provides an optimal balance between solution quality and computational efficiency.
+When the depth counter reaches this threshold, indicating a sustained lack of improvement, the local search terminates and returns the best configuration found.
+Another key innovation in our local search is the incremental fitness calculation.
+Instead of recalculating the entire system’s SRO after each atomic swap, we update only the local regions affected by the swap operation.
+For large systems containing thousands of atoms, this approach reduces the computational overhead by up to 99.6% compared to traditional full recalculation methods.
+The technical details and implementation of this incremental strategy are discussed thoroughly in subsection 4.5.
 4.3 Short-Range Order
-Short-Range Order (SRO) characterizes the local atomic structure in HEAs, describing the tendency of certain atomic species to preferentially cluster compared to a random distribution [37].
-The SRO is typically quantified by the Warren-Cowley parameter [9], denoted as $\alpha_{ij}^{k}$, which measures the degree of ordering between atomic species $i$ and $j$.
-For a given atom $i$, the SRO parameter for its $k$-th coordination shell is defined as:
-$\alpha_{ij}^{k}=1-\frac{P_{ij}^{k}}{c_{j}},$ (2)
-where $P_{ij}^{(k)}$ is the conditional probability of finding a $j$ atom in the $k$-th coordination shell of an $i$ atom, $c_{j}$ is the overall concentration of $j$ atoms in the alloy.
-The SRO parameter $\alpha_{ij}^{k}$ ranges from $[-1,1]$:
-  • 1.
-    $\alpha_{ij}^{k}<0$: Indicates an attraction between elements $i$ and $j$, suggesting a preference to be adjacent.
-  • 2.
-    $\alpha_{ij}^{k}>0$: Indicates a repulsion between elements $i$ and $j$, suggesting a preference to be separated.
-For each atomic pair type $(i,j)$, the SRO parameter for the entire system of $N$ atoms can be expressed as:
-$\langle\alpha_{ij}^{k}\rangle=\frac{1}{N_{i}}\sum_{n=1}^{N_{i}}\left(1-\frac{P_{ij}^{k}}{c_{j}}\right),$ (3)
-where $N_{i}$ is the number of atoms of type $i$.
-To calculate this parameter, the number of $j$ atoms within each coordination shell around every $i$ atom is determined, and the result is normalized by the total concentration of $j$ atoms ($c_{j}$).
-computed SRO values for each pair $(i,j)$ are then averaged over the entire system to obtain a global measure of the SRO.
-4.4 Molecular Dynamics
-Molecular dynamics (MD) simulations were performed using LAMMPS [38, 39, 40], employing the 2-NN MEAM potential for the Fe-Mn-Cr-Co alloy system.
-The simulations encompassed system sizes ranging from 256 to 256,000 atoms, with a timestep of 2 fs in an NPT ensemble.
-The built-in MC method in LAMMPS was utilized to compute the Radial Distribution Function (RDF), providing detailed insight into the atomic arrangement and SRO.
-These simulation parameters were chosen to ensure reproducibility and to facilitate replication in future studies.
-The results of these MD simulations are discussed in detail in section 2, demonstrating the effectiveness of our method.
-4.5 Incremental Calculation of SRO (Fitness)
+Short-range order (SRO) describes how atoms in a multi-component system tend to cluster or repel each other relative to a random distribution [37].
+We quantify SRO using the Warren–Cowley parameter [9], denoted as $\alpha_{ij}^{(k)}$, which characterizes the degree of ordering between atomic species $i$ and $j$ in the $k$-th coordination shell:
+$\alpha_{ij}^{(k)}=1-\frac{P_{ij}^{(k)}}{c_{j}},$ (2)
+where $P_{ij}^{(k)}$ represents the conditional probability of finding a type-$j$ atom in the $k$-th shell around a type-$i$ atom, and $c_{j}$ is the overall concentration of type-$j$ atoms in the system.
+The parameter $\alpha_{ij}^{(k)}$ ranges from $-1$ to $+1$, where negative values indicate preferential attraction between species $i$ and $j$, positive values suggest repulsion, and zero corresponds to a random distribution.
+4.4 Fitness Calculation
+To evaluate how well a given atomic configuration matches a target SRO pattern, we define a fitness function that incorporates the Warren–Cowley parameters:
+$F(X)$ $=\sqrt{\sum_{k=1}^{N_{\mathrm{shells}}}w_{k}\sum_{i=1}^{N_{\mathrm{types}}}\sum_{j=1}^{N_{\mathrm{types}}}\bigl{(}\alpha_{ij}^{(k)}(X)-\alpha_{ij}^{(k),\mathrm{target}}\bigr{)}^{2}}\,,$ (3)
+where $k$ indexes the coordination shells ($1\leq k\leq N_{\mathrm{shells}}$), and $i,j$ enumerate the atomic species ($1\leq i,j\leq N_{\mathrm{types}}$).
+Here, $\alpha_{ij}^{(k)}(X)$ is computed according to Eq.(1) for configuration $X$.
+The shell-specific weights $w_{k}$ and target SRO values $\alpha_{ij}^{(k),\mathrm{target}}$ are user-specified input parameters, allowing flexible control over the optimization process.
+The weights $w_{k}$ can be adjusted to emphasize certain coordination shells, while $\alpha_{ij}^{(k),\mathrm{target}}$ values can be derived from experimental measurements, theoretical calculations, or other simulation results to guide the structure generation toward desired atomic ordering patterns.
+4.5 Incremental Calculation of Fitness
 Figure 6: Comparison of execution times for incremental and non-incremental fitness and total calculations in Monte Carlo search.
@@
 The incremental approach achieves up to a 99.6% reduction in total computation time for larger supercells.
-In the Monte Carlo (MC) search process, the SRO parameter is crucial for evaluating the fitness of each atomic configuration.
-As previously discussed, SRO quantifies the local atomic arrangement, guiding the search toward optimal configurations during each MC iteration.
-In the traditional non-incremental approach, SRO is recalculated from scratch at every MC iteration (Equation 3).
-This recalculation becomes increasingly computationally expensive as system size grows.
-Fig.
-5 illustrates that for larger systems, the SRO calculation can consume more than 99% of the total computational cost, highlighting the necessity for a more efficient method.
-To address this challenge, we implemented an incremental SRO calculation.
-Instead of recalculating the SRO for the entire system at each iteration, we update only the SRO values influenced by the random swap of two atoms, $i$ and $j$.
-This approach localizes the update to the environments of the swapped atoms and their neighbors, enabling a more efficient adjustment of the SRO during each MC iteration.
-For the incremental update, we first compute the change in SRO for the atoms directly involved in the swap. The change in the SRO value for atom $i$ and its neighbors is given by:
-$\Delta\alpha_{ij}^{k}=\left(1-\frac{P_{ij}^{k}\text{(new)}}{c_{j}}\right)-\left(1-\frac{P_{ij}^{k}\text{(old)}}{c_{j}}\right),$ (4)
-[…]
-Benchmark data shown in Fig.
-5 highlights the significant computational burden of non-incremental SRO calculations, which account for 97.4% to 99.6% of the total execution time as system sizes increase from 432 to 3,456 atoms.
-However, the incremental SRO approach substantially alleviates this burden, reducing total execution times to 71.8% for 432 atoms and stabilizing around 65.5% for larger systems.
-These findings underscore the incremental method’s ability to enhance computational efficiency and scalability for large-scale HEA simulations.
-By adopting incremental SRO updates, we maintain the accuracy of fitness evaluations while significantly reducing the computation time for each MC iteration.
-This ensures that the MC search process remains efficient and scalable for large-scale HEA simulations.
+In a non-incremental approach, the Warren–Cowley parameters and resulting fitness are recomputed from scratch at each Monte Carlo step (Equation 2 and Equation 3), which becomes computationally prohibitive for large systems.
+As shown in Fig.
+6, for supercells containing thousands of atoms, these calculations can consume over 99% of the total execution time.
+To address this bottleneck, we implement an incremental fitness update strategy.
+Rather than recalculating all $\alpha_{ij}^{(k)}$ values after each atomic swap, we update only those parameters affected by the exchanged atoms $i$ and $j$.
+The change in the Warren–Cowley parameter for the $k$-th shell is given by:
+$\Delta\alpha_{ij}^{(k)}=\alpha_{ij}^{(k),\mathrm{new}}-\alpha_{ij}^{(k),\mathrm{old}}=\left(1-\frac{P_{ij}^{(k),\mathrm{new}}}{c_{j}}\right)-\left(1-\frac{P_{ij}^{(k),\mathrm{old}}}{c_{j}}\right),$ (4)
+where $P_{ij}^{(k),\mathrm{new}}$ and $P_{ij}^{(k),\mathrm{old}}$ are the probabilities before and after the swap.
+Since only atoms within the local environment of sites $i$ and $j$ are affected, this reduces the computational complexity from $O(N)$ to $O(1)$ per Monte Carlo step.
 4.6 GPU Optimization
-To further accelerate the MC search and enhance scalability, we implemented a parallel GPU-based optimization.
-Each thread processes an independent task, and each thread block manages the complete MC search for a single solution.
-This approach allows the GPU to handle multiple solutions simultaneously, leveraging resources such as registers, shared memory, and low-precision data formats to maximize computational efficiency.
-4.6.1 Parallel Processing on the GPU
-In this setup, each thread performs the following tasks:
+To maximize computational efficiency for large-scale systems, we implement a parallel GPU-accelerated version of the Monte Carlo search.
+The implementation follows a hierarchical parallelization strategy where each GPU thread processes an individual atomic swap operation, while thread blocks manage complete Monte Carlo refinements for distinct candidate solutions.
+This design enables concurrent optimization of multiple configurations while efficiently utilizing GPU resources including registers, shared memory, and mixed-precision arithmetic.
+4.6.1 Parallel Processing Architecture
+The GPU implementation distributes computation across multiple levels:
   • 1.
-    SRO Calculation: Threads compute the SRO for their assigned atomic configurations using low-precision formats (FP16, Uint16) to reduce data footprint and memory access latency, while maintaining sufficient accuracy.
+    Thread-Level Parallelism: Individual threads handle SRO calculations for specific atomic pairs using reduced precision (FP16, Uint16) data types, balancing numerical accuracy with memory efficiency.
   • 2.
-    Lattice Structure Storage: The lattice structure [41] is stored in shared memory to minimize memory access time during each MC iteration.
-This allows all threads within a block to efficiently access and update atomic positions, enhancing computational speed.derived from the SRO values calculated in registers, is updated following each perturbation:
-4.6.2 Iterative Fitness Calculation
-Each thread in the block computes the fitness of the perturbed solution during every MC iteration.
-The fitness, derived from the SRO values calculated in registers, is updated following each perturbation:
+    Block-Level Management: Each thread block maintains a candidate configuration in shared memory, following the lattice structure described in [41].
+This arrangement minimizes global memory access during atomic position updates.
+4.6.2 Incremental Fitness Evaluation
+Following the incremental update strategy detailed in subsection 4.5, each thread computes local fitness changes according to:
 $F_{\text{new}}=F_{\text{prev}}+\Delta F_{\text{perturb}},$ (5)
-where $\Delta F_{\text{perturb}}$ represents the fitness change from the most recent atomic position perturbation. This value is calculated using the incremental SRO update method.
-4.6.3 Fitness Sorting and Selection
-After computing the updated fitness in registers, threads within the block collaborate to sort their fitness values using shared memory.
-The sorting process ensures the rapid identification of the best-performing configuration, which updates the current optimal solution stored in shared memory:
-$F_{\text{best}}=\min\left(F_{\text{new}}\right)$ (6)
-This guarantees that each block retains the best solution found during the current iteration.
-4.6.4 Convergence Check and Termination
-The MC search continues iterating until convergence is achieved.
-Convergence is defined as the point where the fitness of the current best solution shows no significant improvement over a fixed number of steps.
-Specifically, if $\Delta F_{\text{best}}=0$ for a predefined threshold of iterations, the threads within the block terminate the search.
-The final result, including the atomic configuration and the associated fitness, is then written back to global GPU memory for subsequent analysis.
-This GPU-based approach significantly accelerates the MC search by distributing tasks across multiple threads and blocks, enabling the efficient handling of large-scale systems.
+where $\Delta F_{\text{perturb}}$ is calculated from the changes in Warren–Cowley parameters ($\Delta\alpha_{ij}^{(k)}$) affected by the atomic swap.
+This approach maintains consistency with the fitness function defined in Equation 3 while exploiting the locality of updates.
+4.6.3 Parallel Selection Strategy
+After computing fitness changes, threads within each block cooperatively sort the results using shared memory operations. The best configuration update is identified according to:
+$F_{\text{best}}=\min_{\text{updates}}{F_{\text{new}}},$ (6)
+ensuring that each block maintains its optimal configuration throughout the refinement process.
+4.6.4 Convergence Criteria
+The GPU-accelerated search terminates when either:
+  • 1.
+    The fitness value $F_{\text{best}}$ remains unchanged for a number of iterations exceeding the search depth threshold defined in subsection 4.2, or
+  • 2.
+    The maximum number of Monte Carlo steps is reached.
+Upon convergence, the final atomic configuration and its associated fitness value are transferred from GPU to CPU memory for subsequent analysis or further refinement cycles.
+This GPU-optimized implementation achieves significant speedups over CPU-based calculations (see Table 1), enabling the efficient handling of systems containing hundreds of thousands of atoms while maintaining the accuracy of SRO optimization.
+The approach seamlessly integrates with the broader optimization framework described in section 4, providing a scalable solution for large-scale HEA modeling.
+4.7 Molecular Dynamics Simulations
+Molecular dynamics (MD) simulations were performed using LAMMPS [38, 39, 40] for two distinct purposes: generating configurations enriched with SRO features and validating the large-scale structures generated by PyHEA.
+To generate configurations with target SRO properties, MD simulations were conducted on a 500-atom system under specific temperature and pressure conditions.
+An NPT ensemble with a timestep of 2 fs was used, and the simulation ran for 200 ps.
+To accelerate SRO formation, an MC-MD hybrid method was applied, involving random atom swaps every 0.2 ps.
+The final configuration was extracted for subsequent analysis, ensuring alignment with the desired SRO characteristics.
+For validation, MD simulations were carried out on PyHEA-generated configurations containing up to 256,000 atoms.
+These simulations, also performed in an NPT ensemble with a 2 fs timestep, ran for 10 ps.
+RDFs were calculated to evaluate the structural reliability of the large-scale configurations, comparing them against known computational and experimental benchmarks.These complementary simulations ensured that SRO structures were accurately generated in small systems and validated in large-scale configurations, providing a robust framework for modeling high-entropy alloys.
+Overall, these integrated computational techniques—global and local searches, incremental SRO updates, GPU acceleration, and complementary MD—form a robust toolkit for generating and analyzing large-scale HEA models efficiently and accurately.
 5 Acknowledgment
-This work was conducted under the supervision of Professor Lifeng Liu, School of Integrated Circuits,, Peking University.
+This work was conducted under the supervision of Professor Lifeng Liu, School of Integrated Circuits, Peking University.
 The author acknowledges the financial support provided by the National Natural Science Foundation of China (Grant No.
@@
 The table presents computation times (in minutes) and speedup ratios for both body-centered cubic (bcc) and face-centered cubic (fcc) structures of varying sizes (#At) and species numbers (#Sp).
-All simulations maintained consistent parameters: 128 candidate solutions, 10 global iterations, 256 MC tasks (depth 10), and supercell sizes following $S=A\cdot L^{3}$ where $A=2$ (bcc) or $A=4$ (fcc).
-Short-range order was optimized over 3 coordination shells.
-Dashes (–) indicate computationally infeasible cases, while speedup values are shown relative to SCRAPs[ATAT in brackets].
-  Type #Spa #Atb ATAT SCRAPs PyHEA PyHEA Speedupc
-Parallel Parallel GPU CPU GPU
+All simulations maintained consistent parameters: 128 candidate solutions, 10 global iterations, 256 MC tasks, a search depth of 10, and supercell sizes following $S=A\cdot L^{3}$ where $A=2$ (bcc) or $A=4$ (fcc).
+SRO was optimized over 3 coordination shells.
+Dashes (–) indicate computationally infeasible cases, while speedup values are shown relative to SCRAPs[ATAT-mcsqs in brackets].
+  Type #Spa #Atb ATAT- SCRAPs PyHEA PyHEA Speedupc
+mcsqs Parallel Parallel GPU CPU GPU
   bcc 3 54 >1.4kd 0.14 0.14 .03 1.0[10k] 4.67[47k]
@@
   • c
-    Speedup ratios: value vs SCRAPs [value vs ATAT in brackets].
+    Speedup ratios: value vs SCRAPs [value vs ATAT-mcsqs in brackets].
   • d
-    ATAT data from Ref. [23]. All times in minutes.
+    ATAT-mcsqs data from Ref. [23]. All times in minutes.
 Figure 7: Warren–Cowley SRO parameter matrices for a multi-principal-element alloy (Fe, Mn, Cr, Co) at three different temperatures: (a) 1373 K, (b) 873 K, and (c) 873 K.